### v2

# Can Emulating Semantic Translation Help LLMs with Code Translation? A Study Based on Pseudocode

Abstract.
Large language models (LLMs) show great potential in code translation. However, accurate translation remains challenging when using the commonly adopted direct code-to-code translation approach, which converts a program into the target programming language (PL) in a single step. Inspired by the success of incorporating intermediate steps to guide LLMs in resolving challenging tasks, we explore pseudocode-based code translation, which emulates the human semantic translation by first interpreting the program’s intent and logic into pseudocode and then implementing it in the target PL. We find that pseudocode-based translation helps translate programs that direct translation struggles to handle. Nonetheless, the effectiveness, advantages, and limitations of this approach remain underexplored. To bridge this gap, we present an empirical study on pseudocode-based code translation, aiming to investigate its effectiveness in enhancing the direct translation approach, illuminate its effective usage, and identify limitations hindering its potential benefits. By comparing direct and pseudocode-based translation approaches on 9,690 translation tasks across six PLs with five popular LLMs, we demonstrate that pseudocode-based translation can effectively complement direct translation, particularly when translating from flexible to rigid PLs or dealing with low-resource Rust. Based on these findings, we suggest adopting strategies that combine the complementary strengths of both approaches to enhance code translation accuracy. We also reveal the advantages of pseudocode-based translation in disentangling translations of complicated programs and mitigating distractions from detailed implementations in original programs, as well as its limitations due to incorrect, incomplete, or ambiguous pseudocode.
Code Translation, Pseudocode, Semantic Translation, Large Language Model
1. Introduction
Code translation, also known as transpilation, refers to the automatic conversion of a program written in one programming language (PL) to another while preserving its semantics (i.e., program functionality) (Nguyen et al., 2013; Rozière et al., 2020; Pan et al., 2024). With the rapid evolution of PLs and the diverse requirements of software applications, code translation has gained significant attention in both research and industry due to its wide range of practical applications. For example, code translation facilitates the migration of legacy software systems built based on obsolete PLs to modern PLs (Echeverria et al., 2015; Kalia et al., 2021a, b). It also facilitates efficient prototyping and development across multiple PLs for multi-platform software (Liu et al., 2023; Nguyen et al., 2015; Wu et al., 2010). However, achieving accurate automated code translation remains a challenging task due to the inherent differences in syntax and features among various PLs (Pan et al., 2024).
Over the past few decades, various approaches have been proposed to automate code translation and address the associated challenges (Chen et al., 2025; Jiao et al., 2023). Earlier methods primarily relied on statistical or neural machine translation (Nguyen et al., 2013; Chen et al., 2018; Rozière et al., 2020, 2022; Liu et al., 2023), while recent practices start leveraging the powerful large language models (LLMs) as code translators (Pan et al., 2024; Yang et al., 2024). Nevertheless, existing works mainly use a direct code-to-code translation approach, where the original program is taken as input to generate the translated program in a single step. In this step, LLMs may perform multiple tasks implicitly, such as understanding the semantics of the original program and generating the target program in another PL. Studies have shown that such direct code-to-code translation remains challenging, especially across PLs with significant differences in syntax and features (Pan et al., 2024; Macedo et al., 2025).
Decomposing challenging tasks into intermediate steps has proven effective in guiding LLMs to emulate successful human workflows in various coding tasks (e.g., splitting planning and implementation for code generation (Dong et al., 2024; Li et al., 2025) and fault localization and patch generation for debugging (Xia et al., 2024)). Notably, we observe that code translation can also benefit from explicitly implementing intermediate steps that emulate human practices in translation. Specifically, human translators often perform semantic translation (Newmark, 1981) based on the meaning of the given text, first distilling the intent into a language-agnostic meaning and then rendering it in the target language (Newmark, 1981). Meanwhile, a recent study demonstrates that pseudocode, as a PL-agnostic representation of code intent and logic widely used in textbooks and research papers, can effectively guide code generation across PLs (Wu et al., 2025). Based on these insights, we further recognize that emulating semantic translation for code based on pseudocode can be a solution for the tasks that direct translation struggles with. For example, when translating a simple C++ program (Figure 1(a)) to Rust, all five studied LLMs (e.g., Qwen2.5-Coder-32B-Instruct) failed to produce an accurate translation via a one-step direct translation (Figure 1(b)). In comparison, following a clear pseudocode (Figure 1(c)) summarized from the original C++ program, these LLMs can successfully implement the original program’s functionality in Rust (Figure 1(d)). However, introducing explicit intermediate steps also has limitations. For example, it may introduce noise or suffer from semantic loss during information transmission (Liu et al., [n. d.]), which may conversely hinder accurate code translation. In general, the effectiveness, advantages, and limitations of such pseudocode-based code translation compared to the commonly adopted direct translation remain unclear.
In this work, we conduct the first empirical study on pseudocode-based code translation to bridge the gap in understanding this under-explored approach. Specifically, we assessed the performance of five translation strategies: the widely-adopted direct translation approach, two pseudocode-based strategies, and two hybrid strategies that combine direct and pseudocode-based translation. Through four research questions (RQs), we investigate the overall effectiveness of pseudocode-based code translation compared with the widely-adopted direct translation (RQ1), the helpfulness of pseudocode for varying PL pairs (RQ2), the advantages of pseudocode as an intermediary compared to a concrete PL, which also proved helpful in code translation (Macedo et al., 2025) (RQ3), and how the quality of pseudocode affects translation outcomes (RQ4). We also examine the successes and failures of pseudocode-based code translation via case studies and discuss future directions for further harnessing its potential benefits. The experiment is conducted with six popular PLs on solution programs of 323 LeetCode problems in three difficulty levels in LiveCodeBench (Jain et al., 2025), resulting in 9,690 translation tasks for each of five studied LLMs.
Our experimental results reveal several interesting findings and actionable insights. Specifically, we identify the complementary role of pseudocode-based code translation in effectively handling code translation tasks that direct translation struggles with. By combining the results of direct translation and pseudocode-based translation, the studied LLMs achieve average improvements in pass@10 rates over direct translation alone of 4.10%, 7.44%, and 13.75% on easy-, medium-, and hard-level problems, respectively. We also find that pseudocode-based translation can benefit varying source-target PL pairs, with more substantial helpfulness for translation from flexible PLs (e.g., Python) to more rigid PLs (e.g., Go) or involving the low-resource PL Rust. These findings suggest the adoption of a hybrid strategy combining direct and pseudocode-based translation results with further test-based selection to leverage the complementary advantages of both approaches, in particular when translating from flexible to rigid PLs or dealing with low-resource PLs. Moreover, with higher-quality pseudocode, we observed that the studied LLMs show the potential to achieve pass rates of 0.9646–0.9835, 0.8861–0.9512, and 0.6747–0.8286 on three-level tasks, respectively. The identified bottlenecks of pseudocode-based code translation stem from both code understanding and generation capabilities of LLMs. Our case studies also highlight the advantages of pseudocode-based translation in disentangling understanding and generation on complicated programs and mitigating distractions from the PL-specific details in original programs, as well as its limitations caused by incorrect, incomplete, or ambiguous pseudocode. These findings further illuminate the effective use of pseudocode-based translation and inspire future research on enhancing code translation through pseudocode.
Figure 1. An Example of Correct C++-to-Rust Translation by Qwen2.5-Coder-32B-Instruct based on Pseudocode (The other four studied LLMs show similar symptoms in this case.)
To summarize, this work makes the following contributions:
  • We conduct the first empirical study on pseudocode-based code translation, exploring the effectiveness, advantages, and limitations of explicit semantic code translation via pseudocode in improving code translation accuracy by emulating human translation practices.
  • We compare four pseudocode-based translation strategies with direct translation across 9,690 translation tasks based on programs in six popular PLs. With five popular LLMs as the code translator, we systematically investigate the effectiveness of pseudocode-based code translation in enhancing direct translation, its helpfulness for varying PL pairs, and its advantages and limitations, aiming to provide insights for its effective use and inspire future enhancement.
  • Our study reveals interesting findings, including the complementary role of pseudocode-based translation in enhancing direct code translation, particularly for flexible-to-rigid PL translation and for low-resource Rust, and the potential and bottlenecks of pseudocode-based code translation. These findings yield actionable insights, including the adoption of a hybrid strategy combining direct and pseudocode-based translation results to enhance code translation accuracy, as well as the space in pseudocode quality to further harness the benefits of the approach.
The rest of the paper is organized as follows. Section 2 introduces the background of code translation and motivating examples. Section 3 presents the design of our empirical study, including the research questions, studied translation strategies, translation tasks, LLMs, and evaluation metrics. Section 4 analyzes the experimental results to answer our research questions in detail and discusses the advantages and limitations of pseudocode-based code translation through case studies. Section 6 discusses related work. Finally, Section 7 concludes the paper.
2. Background and Motivation
2.1. Code Translation
Code translation refers to the process of converting programs written in one programming language (PL) into another PL while preserving code semantics (Nguyen et al., 2013; Rozière et al., 2020; Pan et al., 2024). Formally, given an original program $S=\langle s_{1},s_{2},\dots,s_{n}\rangle$ written in the source PL $L_{s}$, code translation aims to produce a translated program $T=\langle t_{1},t_{2},\dots,t_{m}\rangle$ in the target PL $L_{t}$, such that $T$ performs the same functionality as $S$, where $s_{i}$ and $t_{j}$ represent statements in $L_{s}$ and $L_{t}$, respectively. For example, we can translate the C++ program in Figure 1(a) and the Python program in Figure 2(a) to their semantically equivalent Rust program in Figure 1(d) and Java program in Figure 2(d), respectively. In code translation, there may not always be equivalent constructs, statements, and APIs in the target PL for those in the source PL, and vice versa, which makes code translation challenging.
In the past decade, automated code translation approaches have developed from traditional rule-based and statistical machine translation (Nguyen et al., 2013) to neural-model-based methods to automatically learn diverse and complex translation patterns across PLs (Chen et al., 2018; Rozière et al., 2020, 2022; Liu et al., 2023; Jiao et al., 2023). Recently, LLMs pre-trained on massive code corpora demonstrate superior capabilities in code translation (Yang et al., 2024; Ibrahimzada et al., 2025; Pan et al., 2024). These works mainly focus on a code-to-code translation approach, i.e., directly converting the original program in $L_{s}$ to the translated program in $L_{t}$ in an integrated step. However, LLMs may struggle to precisely capture and replicate the semantics of the original program in another PL in a single step, leading to incorrect translations (Pan et al., 2024). To mitigate the gap between source and target PLs, Macedo et al. (2025) explored transitive translation via an intermediate PL, but they still focus on code-to-code translation and fail to handle certain translation tasks. In this study, we explore an alternative code translation approach by emulating human semantic translation in natural language translation (Newmark, 1981), which explicitly conducts code understanding by generating pseudocode as an intermediate step to interpret code intent and logic to facilitate translation.
In code translation, the existence of original programs provides a reference to validate the translated programs. Unlike code generation tasks that rely on manually defined test cases or human evaluation to assess the generated code (Chen et al., 2021; Austin et al., 2021), code translation allows automated correctness assessment by comparing the outputs of the translated and original programs on the same test cases. These enable multiple attempts to be a feasible and meaningful practice to enhance code translation accuracy (Macedo et al., 2025; Zheng et al., 2023). In this study, we explore the translation setup of 10 attempts, which is a typical cost-effective setup that balances cost and gain (Macedo et al., 2025; Wu et al., 2025). It also enables the investigation of combining results from both direct and pseudocode-based code translation (Section 3.3).
2.2. Motivating Examples
Figure 2. An Example of Correct Python-to-Java Translation by DeepSeek-Coder-V2-Lite based on Pseudocode
To guide LLMs to solve complicated coding tasks, researchers often decompose a complex task into multiple simpler sub-tasks following human experience (e.g., problem solving/planning and implementation for code generation (Dong et al., 2024; Li et al., 2025; Wu et al., 2025)). Inspired by this, we explore whether decomposing code translation into sub-tasks following human practices can also facilitate LLMs to produce more accurate translations by emulating the successful human workflow.
Recalling the practices in natural language translation, humans typically employ two approaches. When expressions or sentence structures in the source and target languages are sufficiently similar, a literal word-by-word translation is often adequate. However, when literal translation is challenged by significant differences in grammar, idiomatic usage, or context between the source and target languages, humans tend to adopt semantic translation (Newmark, 1981), where they first interpret the underlying meaning of the source sentence and then render it in the target language in the manner best to convey the intended message.
Interestingly, we observe that semantic interpretation also benefits LLMs in code translation. Specifically, we prompt LLMs to explicitly conduct semantic translation via pseudocode-based transitive translation, i.e., first producing a semantic interpretation (pseudocode) for code intent and logic of the original program, and then generating the translated program based on pseudocode. We adopt pseudocode to represent semantics since it is a widely adopted form for describing code intent and logic of algorithms in textbooks and research papers. It also proves effective in guiding LLMs to replicate the code intent and logic in varying PLs and can be synthesized from programs (Wu et al., 2025).
Figure 2 shows an example where pseudocode enables DeepSeek-Coder-V2-Lite (abbr. DSCoder) to correctly translate a Python program that cannot be accurately transformed through direct translation. Specifically, when translating the Python program in Figure 2(a) into Java directly, DSCoder attempts to mimic the Python implementation by using a for loop. Although it recognizes that Java does not support lambda expressions within for loops, the resulting translation (Figure 2(b)) employs in-place manipulation that is structurally similar but semantically inconsistent with the original Python program. In contrast, when DSCoder first abstracts the intent of the Python program into pseudocode (Figure 2(c)) and then generates Java code from this pseudocode, it produces a correct implementation by updating the array as a whole via a temporary array (Figure 2(d)).
Moreover, we observe that the semantic code translation via pseudocode is broadly helpful for LLMs in various families and sizes. As shown in Figure 1, when translating a simple C++ solution of an easy-level LeetCode programming problem into Rust, the powerful Qwen2.5-Coder-32B-Instruct (abbr. Qwen32B) fails to produce a correct translation via direct translation, by misusing the log2 API in Rust as shown in Figure 1(b). However, when guided by the corresponding pseudocode of the original program (Figure 1(c)), Qwen32B correctly leverages the leading_zeros API to achieve the intended functionality as shown in Figure 1(d). Notably, all five studied LLMs (listed in Table 1) suffer from similar issues, and all of them except DSCoder generate a correct implementation when provided with pseudocode summarized by themselves as an intermediate step.
These motivating examples suggest that incorporating pseudocode as a semantic representation for transitive translation can promote accurate code translations, mirroring the benefits of humans using semantic translation for natural languages. This is reasonable as the approach guides LLMs to understand code intent and logic and then perform code generation, decomposing a complex task into simpler sub-tasks, and LLMs are skillful in both tasks (Geng et al., 2024; Jain et al., 2025). This enables a new access to achieving code translation, which may bypass LLMs’ struggle with direct translation.
Despite these benefits, the proper use of pseudocode in code translation remains underexplored, e.g., how to effectively translate code via pseudocode, how effective it is when handling distinct source-target PL pairs, and what strengths it entails. In addition, pseudocode-based code translation may also face challenges, e.g., the potential error propagation from incorrect pseudocode and the inherent ambiguity in natural language. The limitations it entails and potential remedies to overcome the limitations also remain unclear. To bridge these gaps, we conduct a systematic investigation of pseudocode-guided code translation. Our study aims to provide deeper insights into when and how pseudocode can effectively enhance code translation.
3. Study Design
3.1. Research Questions
In this study, we investigate the effectiveness of pseudocode-based code translation by studying the following four research questions (RQs).
RQ1: How effective is pseudocode-based code translation?
This RQ aims to show a general picture of the effectiveness of semantic translation for code via pseudocode. We also investigate how such pseudocode-based code translation can work with the widely adopted direct code translation, e.g., fully replacing direct translation or complementing it by adopting both strategies simultaneously.
RQ2: How effective is pseudocode-based semantic translation for different pairs of source and target programming languages?
Existing studies show that the code translation accuracy of LLMs varies across different pairs of source and target PLs, indicating that distinct PL pairs may pose varying challenges for LLMs (Pan et al., 2024; Macedo et al., 2025). Pseudocode may also have varying helpfulness for different source and target PL pairs. Therefore, this RQ verifies whether pseudocode-based translation is generally helpful for all PL pairs. It also investigates whether certain PL pairs can benefit more from pseudocode-based translation, suggesting scenarios where pseudocode-based translation is highly recommended.
RQ3: Is pseudocode a more effective intermediary compared to a specific programming language?
Macedo et al. (2025) demonstrates that a specific PL with close syntactic and semantic similarities to the source and target PLs may act as an effective transitive intermediary to benefit LLMs in code translation. For example, translating Python code to Java via Rust is found to lead to better accuracy than direct translation. In this RQ, we aim to understand whether pseudocode can serve as a more effective intermediary than a specific transitive PL.
RQ4: How does the quality of pseudocode affect the effectiveness of pseudocode-based code translation?
LLMs may not accurately interpret and record the intent of the input program, resulting in low-quality pseudocode that may conversely hinder LLMs from generating correct code translation. This RQ investigates the code translation performance based on high-quality pseudocode, aiming to reveal the potential of pseudocode-based code translation and identify the bottlenecks of studied LLMs to harness the potential effectiveness.
3.2. Experimental Translation Subjects
3.2.1. Programming Languages
In this study, we evaluate code translation performance across six widely studied PLs, i.e., C++, Python, Java, JavaScript, Go, and Rust. The choice of these PLs follows prior work, where translation among the first four PLs has been extensively explored (Rozière et al., 2022; Jiao et al., 2023; Yang et al., 2024), while recent studies have highlighted the emerging PLs such as Go and Rust (Pan et al., 2024; Macedo et al., 2025).
3.2.2. Code Translation Tasks
We prompt LLMs to translate the solution programs of 323 LeetCode problems in LiveCodeBench (Jain et al., 2025). We choose to translate such programs since they involve common implementations widely used in practical development and become a representative benchmark to evaluate the coding ability of LLMs (Hui et al., 2024; Team, 2024). Actually, solution programs of LeetCode problems have been widely adopted as the subjects in code translation studies since they formulate easy-to-collect multilingual parallel programs (Wang et al., 2023a; Ahmad et al., 2023). Moreover, LiveCodeBench provides carefully validated and representative tests, facilitating the reliable validation of the translation results. Additionally, the LiveCodeBench programming problems are released after May 2023 (Jain et al., 2025), allowing the evaluation to be less subject to the data contamination issue compared to the conventional code translation benchmarks like CodeNet (Puri et al., 2021), TransCoder (Rozière et al., 2020), and AVATAR (Ahmad et al., 2023) established in 2020–2021.
To support the investigation of code translation among diverse PLs, we follow existing practices (Jain et al., 2025; Wu et al., 2025) to collect correct user-written solution programs that can pass all the tests on LiveCodeBench, for all six PLs. These solution programs are used as the source programs to translate. Among all 381 LeetCode problems from LiveCodeBench, we excluded 16 problems with incorrect tests identified by Wu et al. (2025). We also excluded another 42 problems lacking solutions in at least one of the six PLs on LeetCode. This results in 116 easy problems, 160 medium problems, and 47 hard problems. Note that, for each problem, LiveCodeBench does not provide corresponding solution programs and only supports the evaluation of Python solutions. We collect the solution programs in six PLs and extend the evaluation scaffold to all six studied PLs via execution-based validation, following Wu et al. (2025) (who extended the support of C++ and Rust). We open-source these materials to facilitate future research on code translation (Artifact of this paper, [n. d.]).
[…]
3.3. Experimental Strategies for Code Translation
In this study, we investigate the effectiveness of five prompting strategies for code translation. These include one strategy based on the commonly adopted direct translation approach, two pseudocode-based translation strategies, and two hybrid translation strategies.
  1. (1)
    Direct Translation (abbr. $\mathcal{D}$): This strategy prompts LLMs to translate the programs in the source PL to the target PL. The strategy is widely adopted by existing LLM-based code translation studies (Yang et al., 2024; Pan et al., 2024). We follow these studies to design the prompt as shown in Figure 3(a).
  2. (2)
    Transitive Translation via Pseudocode (abbr. $\mathcal{P}$): This strategy prompts LLMs to explicitly conduct semantic translation using two transitive steps: (i) generate a piece of pseudocode based on the original program following the prompt from Wu et al. (2025); and (ii) implement the pseudocode into a translated program in the target PL following the prompt in Figure 3(b). As introduced in Section 2.2, this strategy is inspired by human practices of translation and the findings that pseudocode can encode the PL-agnostic code intent and logic of a program (Wu et al., 2025), serving as an appropriate semantic representation for emulating semantic translation. In our study, pseudocode and translated programs are generated by the same LLM (referred to as “translator LLM”). The prompts are designed following the practices of Wu et al. (2025).
  3. (3)
    Transitive Translation via Pseudocode with Original Program as Context (abbr. $\mathcal{PC}$): This strategy extends the above pseudocode-based strategy to include more implementation details (i.e., original program) as the context to guide code translation. Specifically, considering that pseudocode abstracts implementation details of the original program, it may miss some helpful information for LLMs to implement the original program’s intent and logic in the target PL. Thus, we include the original program in the source PL as an implementation example of the pseudocode for LLMs’ reference when generating the target code through explicit semantic translation, using the prompt shown in Figure 3(c).
As mentioned in Section 2.1, we generate ten translations for each original program. When using the above strategies, we repeatedly generate ten translations using their respective prompts.
We also examine the effectiveness of combining direct translation with pseudocode-based translation. Specifically, we design two straightforward hybrid strategies by evenly mixing translation results of direct translation and transitive translation via pseudocode, while keeping the same total attempts as the above strategies (i.e., 10 attempts as mentioned in Section 2.1, with five attempts for each approach). The two hybrid strategies are defined as follows:
  1. (4)
    Hybrid translation $\mathcal{D}\&\mathcal{P}$: This strategy mixes five translation results from Direct Translation ($\mathcal{D}$) and another five translated results from Transitive Translation via Pseudocode ($\mathcal{P}$).
  2. (5)
    Hybrid translation $\mathcal{D}\&\mathcal{PC}$: This strategy mixes five translation results from Direct Translation ($\mathcal{D}$) and another five translated results from Transitive Translation via Pseudocode with Original Program as Context ($\mathcal{PC}$).
We include these hybrid strategies to study whether combining direct translation and pseudocode-based translation can leverage their complementary strengths across different translation tasks.
(a) Prompt for Direct Translation
(b) Prompt for Generating Code from Pseudocode
(c) Prompt for Generating Code from Pseudocode with Source Program as Reference Implementation
Figure 3. Prompts Used in the Study (The examples are instantiated with the translation tasks in Figure 1).
3.4. Experimental Setups
Metric: We report the pass@10 rate of the translated programs (i.e., programs in the target PL) on the test cases provided by LiveCodeBench to reflect the computational accuracy of the translation results. The metric measures the ratio of tasks on which the translator LLM successfully generates a correct translation in at least one out of ten attempts. Compared with text-based metrics like BLEU and Exact Match rate, the execution-based pass@10 metric can measure the semantic equivalence of the translation results based on the high-quality tests provided in LiveCodeBench. The metric is widely adopted in code generation (Jain et al., 2025; Wu et al., 2025) and translation studies (Macedo et al., 2025). Note that we use pass@10 rather than using pass@1 for one attempt, because we aim to investigate the maximum potential of the studied strategies and LLMs in code translation, considering the higher translation accuracy achieved with ten attempts than one single attempt demonstrated by existing study (Macedo et al., 2025), as well as the practical feasibility of filtering out the incorrect translations based on the source program’s behavior as discussed in Section 2.1. The result of pass@1 shows similar trends.
Studied LLMs: In this study, we investigate five popular LLMs from different famous LLM families deployable on our machine. They are Qwen2.5-Coder-32B-Instruct (abbr. Qwen32B) (Hui et al., 2024), Qwen2.5-Coder-7B-Instruct (abbr. Qwen7B) (Hui et al., 2024), Phi-4 (Abdin et al., 2024), GPT-4o-mini (OpenAI, [n. d.]), and DeepSeek-Coder-V2-Lite-Instruct (abbr. DSCoder) (DeepSeek-AI, 2024). They include both general-purpose LLMs and coding LLMs, and both open-sourced and commercial LLMs from different families. The detailed information of these LLMs is listed in Table 1.
Table 1. Information of LLMs Involved in the Study
LLM Name Family Type Parameter Size Knowledge Cutoff
Qwen2.5-Coder-32B-Instruct*
(abbr. Qwen32B)
Qwen Open Source 32.8B Mar 2024 (Knowledge Cutoff Information of Qwen32B, [n. d.])
Phi-4 Microsoft Open Source 14.7B May 2024 (Knowledge Cutoff Information of Phi-4, [n. d.])
GPT-4o-mini OpenAI Commercial (Unknown) Oct 2023 (Knowledge Cutoff Information of GPT-4o-mini, [n. d.])
Qwen2.5-Coder-7B-Instruct
(abbr. Qwen7B)
Qwen Open Source 7.62B Mar 2024 (Knowledge Cutoff Information of Qwen7B, [n. d.])
DeepSeek-Coder-V2-Lite
-Instruct (abbr. DSCoder)
DeepSeek Open Source 15.7B Nov 2023 (Knowledge Cutoff Information of DSCoder, [n. d.])
*: We run Qwen2.5-Coder-32B-Instruct-GPTQ-Int4 on with NVIDIA RTX4090 GPU. The quantized GPTQ-Int4 version is found to show comparable coding ability as the original bf16 version while taking less GPU memory (Wu et al., 2025).
LLM Configuration: We run the four open-source LLMs on our server and access the commercial GPT-4o-mini via OpenAI API. To balance the diversity and reliability, we adopt the temperature of 0.2 following the practice in code generation studies (Jain et al., 2025; Wu et al., 2025; Xu et al., 2024). The maximum output tokens are set to be 3000 (i.e., 1.5x maximum input code length) to accommodate normal code translation outputs. The other configurations (e.g., top_p and penalty) are kept as default values. To mitigate randomness, we repeat the experiments three times and report the average results.
Experimental Environment: We run experiments on our server with Ubuntu 22.04 OS. The server is equipped with NVIDIA RTX4090 GPUs to deploy the open-source LLMs, as well as an AMD Ryzen Threadripper PRO 3995WX 64-Core CPU to support the parallel evaluation of massive generated translation results.
4. Results and Analysis
4.1. RQ1: Overall Effectiveness of Pseudocode-based Transitive Code Translation
In RQ1, we investigate the overall effectiveness of pseudocode-based code translation. Specifically, we explore (1) how pseudocode-based code translation can help enhance the performance of the current widely-adopted direct translation strategy; and (2) whether pseudocode generated by the translator LLM can effectively represent the code intent or logic of the original program under translation and promote accurate code translation results. We investigate these by comparing the performance of five translation strategies introduced in Section 3.3. We take the widely adopted direct translation strategy ($\mathcal{D}$) as the baseline and analyze the performance improvement of the other four strategies. Table 2 shows the pass@10 rate of each LLM using different strategies.
Table 2. Pass@10 Rate of Code Translation with Different Translation Strategies
LLM Strategy Easy Tasks Medium Tasks Hard Tasks
Qwen32B $\mathcal{D}$  (Direct Translation) 0.9480 0.8772 0.7059
$\mathcal{P}$  (viaPseudocode) 0.9442 (-0.40%) 0.8736 (-0.41%) 0.6768 (-4.12%)
$\mathcal{PC}$  (viaPseudocode w/ program as context) 0.9648 (+1.77%) 0.9001 (+2.61%) 0.7553 (+7.00%)
$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudo.]) 0.9773 (+3.09%) 0.9403 (+7.19%) 0.7948 (+12.59%)
$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudo. w/ prog.]) 0.9773 (+3.09%) 0.9327 (+6.33%) 0.7988 (+13.16%)
Phi-4 $\mathcal{D}$  (Direct Translation) 0.9276 0.8313 0.6281
$\mathcal{P}$  (viaPseudocode) 0.8782 (-5.33%) 0.7744 (-6.84%) 0.5832 (-7.15%)
$\mathcal{PC}$  (viaPseudocode w/ program as context) 0.9254 (-0.24%) 0.8347 (+0.41%) 0.6463 (+2.90%)
$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudo.]) 0.9646 (+3.99%) 0.8947 (+7.63%) 0.7083 (+12.77%)
$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudo. w/ prog.]) 0.9610 (+3.60%) 0.8903 (+7.10%) 0.7038 (+12.05%)
GPT-4o-mini $\mathcal{D}$  (Direct Translation) 0.9279 0.8552 0.6667
$\mathcal{P}$  (viaPseudocode) 0.8647 (-6.81%) 0.7219 (-15.59%) 0.5000 (-25.00%)
$\mathcal{PC}$  (viaPseudocode w/ program as context) 0.9239 (-0.43%) 0.8635 (+0.97%) 0.6837 (+2.55%)
$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudo.]) 0.9658 (+4.08%) 0.9054 (+5.87%) 0.7404 (+11.05%)
$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudo. w/ prog.]) 0.9621 (+3.69%) 0.9144 (+6.92%) 0.7617 (+14.25%)
Qwen7B $\mathcal{D}$  (Direct Translation) 0.9073 0.7928 0.5400
$\mathcal{P}$  (viaPseudocode) 0.7735 (-14.75%) 0.5928 (-25.23%) 0.3286 (-39.15%)
$\mathcal{PC}$  (viaPseudocode w/ program as context) 0.8989 (-0.93%) 0.7929 (+0.01%) 0.5480 (+1.48%)
$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudo.]) 0.9336 (+2.90%) 0.8395 (+5.89%) 0.5849 (+8.31%)
$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudo. w/ prog.]) 0.9420 (+3.82%) 0.8568 (+8.07%) 0.6154 (+13.96%)
DSCoder $\mathcal{D}$  (Direct Translation) 0.8964 0.7876 0.5586
$\mathcal{P}$  (viaPseudocode) 0.8562 (-4.48%) 0.7335 (-6.87%) 0.5069 (-9.26%)
$\mathcal{PC}$  (viaPseudocode w/ program as context) 0.9218 (+2.83%) 0.7875 (-0.01%) 0.5530 (-1.00%)
$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudo.]) 0.9482 (+5.78%) 0.8569 (+8.80%) 0.6442 (+15.32%)
$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudo. w/ prog.]) 0.9537 (+6.39%) 0.8581 (+8.95%) 0.6459 (+15.63%)
(Average) $\mathcal{D}$  (Direct Translation) 0.9214 0.8288 0.6199
$\mathcal{P}$  (viaPseudocode) 0.8634 (-6.30%) 0.7392 (-10.81%) 0.5191 (-16.26%)
$\mathcal{PC}$  (viaPseudocode w/ program as context) 0.9270 (+0.60%) 0.8357 (+0.83%) 0.6373 (+2.81%)
$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudo.]) 0.9579 (+3.96%) 0.8874 (+7.06%) 0.6945 (+12.04%)
$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudo. w/ prog.]) 0.9592 (+4.10%) 0.8905 (+7.44%) 0.7051 (+13.75%)
Effectiveness in Enhancing Direct Translation. The results show that the two hybrid strategies (i.e., $\mathcal{D}\&\mathcal{P}$ and $\mathcal{D}\&\mathcal{PC}$) that combine results from pseudocode-based and direct translation effectively improved the code translation accuracy achieved by the baseline ($\mathcal{D}$). Specifically, $\mathcal{D}\&\mathcal{P}$ improved the pass@10 rate by 3.96%, 7.06%, and 12.04% on average on three difficulty levels, respectively. $\mathcal{D}\&\mathcal{PC}$ led to more significant improvements of 4.10%, 7.44%, and 13.75% on three levels on average, respectively. The results demonstrate that pseudocode-based code translation enables LLMs to correctly translate many programs that cannot be correctly handled by direct translation alone. Thus, given the same number of attempts, it is beneficial to adopt both direct translation and pseudocode-based translation instead of consistently giving chances to either strategy. The results indicate that pseudocode-based code translation can effectively complement the widely adopted direct translation to enhance code translation accuracy.
Meanwhile, we observed that the two pure pseudocode-based translation strategies (i.e., $\mathcal{P}$ and $\mathcal{PC}$) only produce comparable or even worse performance than direct translation ($\mathcal{D}$). Specifically, $\mathcal{PC}$ increased pass@10 of $\mathcal{D}$ by only 0.60%, 0.83%, and 2.81% on three levels on average; $\mathcal{P}$ even led to drops in pass@10 rate by 6.30%, 10.81%, and 16.26%. The results demonstrate that when conducting semantic translation via pseudocode, LLMs also failed to translate a few programs that could be correctly translated by direct translation, indicating that direct translation and pseudocode-based translation are complementary at the current stage. We discuss potential causes of this underperformance via case studies in Section 4.5; we also explore the potential of pure pseudocode-based code translation based on higher-quality pseudocode in RQ4 (Section 4.4).
Difficulty-wise. We also notice that the improvement from hybrid strategies is more significant for harder tasks. Specifically, the improvement in pass@10 rate brought by $\mathcal{D}\&\mathcal{P}$ and $\mathcal{D}\&\mathcal{PC}$ over $\mathcal{D}$ increases by 3.96% and 4.10% on easy-level tasks, while 12.04% and 13.75% on hard-level tasks on average, respectively. The results indicate that pseudocode-based code translation is notably helpful to complement direct translation on relatively complicated programs.
Message 1: Pseudocode-based code translation can effectively complement the widely adopted direct translation strategy to translate programs that cannot be correctly handled by direct translation alone, with a more significant improvement for relatively complicated programs. We recommend practitioners adopt hybrid strategies that combine the strengths of direct translation and pseudocode-based translation to enhance code translation accuracy, by collecting translated programs from both approaches and identifying the final result based on tests.
Usefulness of Original Programs. The comparison between the pure pseudocode-based strategies with and without the original program (i.e., $\mathcal{P}$ vs. $\mathcal{PC}$) further reveals the need to include the original program in pseudocode-based code translation. Specifically, the pure pseudocode-based strategy with the original program as context ($\mathcal{PC}$) achieved an average pass@10 rate of 0.9270, 0.8357, and 0.6373 on the three levels’ solution programs, respectively, while the strategy without the original program ($\mathcal{P}$) only achieved pass@10 rates of 0.8634, 0.7392, and 0.5191 accordingly. The results demonstrate that the pseudocode generated by the translator LLM missed or misinterpreted details necessary for the correct generation of programs in the target PL in some cases. We investigate the concrete symptoms of this issue via case studies and discuss them in Section 4.5.
Meanwhile, the helpfulness of including original programs becomes less obvious for the hybrid strategies (i.e., $\mathcal{D}\&\mathcal{P}$ vs. $\mathcal{D}\&\mathcal{PC}$). Specifically, the pass@10 rates of $\mathcal{D}\&\mathcal{P}$ and $\mathcal{D}\&\mathcal{PC}$ are generally comparable (with less than 0.01 difference in pass@10) in most comparisons. This indicates that the original program is less necessary to supplement pseudocode when using hybrid strategies, where direct translation may already provide access to referencing details in the original program for cases that highly rely on such details. The exceptional cases include GPT-4o-mini on hard-level tasks and Qwen7B on medium-level and hard-level tasks, where the hybrid strategy with the original program as context ($\mathcal{D}\&\mathcal{PC}$) achieved higher pass@10 rates. Therefore, it may still be beneficial to include the original program as a backup reference implementation of pseudocode when translating relatively complicated programs.
Message 2: The pseudocode generated by translator LLMs may miss details necessary for generating semantically equivalent programs in the target PL. Including the original program as context is generally suggested to harness the benefits of pseudocode-based code translation, especially for complicated programs.
4.2. RQ2: Effectiveness Across Different Source and Target Programming Languages
Differences in syntax and semantics among PLs may cause varying translation difficulty across PL pairs (Pan et al., 2024; Macedo et al., 2025). In this RQ, we thus further study the effectiveness of the pseudocode-based translation strategies on different PL pairs. The comparison aims to (1) verify whether the helpfulness of pseudocode-based translation identified in RQ1 persists across different source and target PLs, and (2) explore if the pseudocode-based translation leads to significant improvements for certain PL pairs. The empirical findings are intended to guide developers in adopting pseudocode-based approaches for code translation involving different PLs. In this RQ, we focus on the improvement of the hybrid translation strategy with code ($\mathcal{D}\&\mathcal{PC}$), which exhibited generally optimal performance in RQ1, relative to the conventional direct translation strategy ($\mathcal{D}$).
(a) Qwen-32B (Easy Tasks)
(b) Qwen-32B (Med Tasks)
(c) Qwen-32B (Hard Tasks)
(d) Phi-4 (Easy Tasks)
(e) Phi-4 (Med Tasks)
(f) Phi-4 (Hard Tasks)
(g) GPT-4o-mini (Easy Tasks)
(h) GPT-4o-mini (Med Tasks)
(i) GPT-4o-mini (Hard Tasks)
(j) Qwen-7B (Easy Tasks)
(k) Qwen-7B (Med Tasks)
(l) Qwen-7B (Hard Tasks)
(m) DSCoder (Easy Tasks)
(n) DSCoder (Med Tasks)
(o) DSCoder (Hard Tasks)
Figure 4. Relative Improvement in Pass@10 Using “Direct + viaPseudocode (w/ code)” Strategy Compared to “Direct” Strategy for Code Translation. (The greener, the more improvement.)
Overall Effectiveness. Figure 4 presents the relative improvement in pass@10 rates of the hybrid strategy $\mathcal{D}\&\mathcal{PC}$ over the baseline $\mathcal{D}$ across different PL pairs. Each row/column in a heatmap corresponds to a specific source/target PL. In general, we observed that the translation accuracy between almost all PL pairs improves (as indicated by positive values in figures) when using the $\mathcal{D}\&\mathcal{PC}$ strategy compared to $\mathcal{D}$. Across all 450 PL-pair-model-difficulty combinations, 147 (32.7%) combinations achieve even more than 10% improvement. The improvement is consistently observed across different LLMs and task difficulties, with more significant improvement on harder tasks and weaker LLMs, as observed in RQ1. The results confirm the generalizable benefits of adopting pseudocode-based translation among different PL pairs.
Comparison Among PL Pairs. As shown by the different colors in heatmaps within Figure 4, the improvement varies across different PL pairs, which may be due to the differences across PL pairs and the varying training corpus of different PLs. Nevertheless, we observe consistent trends across different LLMs and task difficulties. Specifically, we first noticed that the improvement of code translation from flexible and lightweight PLs (e.g., Python and JavaScript) to stricter and more complex PLs (e.g., Rust and Go) is usually more significant than the improvement in the opposite direction, as shown by the deeper green colors near the right-bottom corner of each heatmap. For example, when conducting Python-to-Rust translation and JavaScript-to-Go translation on medium-level tasks, the improvement on Qwen32B is 13.64% and 8.96%, respectively; in comparison, it is only 5.54% and 4.05% in the opposite direction, respectively. We conjecture this is because transitive translation based on pseudocode helps guide LLMs to generate code fitting the requirements enforced in stricter and more complex PLs (e.g., Python and JavaScript are flexible with dynamic typing and rich syntax sugar, while C++ and Java require static type declaration and more rigid syntax, and Go and Rust further enforce more rules on variable usage and memory management).
In addition, we observed that code translation to Rust (rightmost column in heatmaps) generally gains significant improvement, regardless of the source PL. We conjecture this may also result from the relatively lower resource of the Rust training corpus (Baltaji et al., 2025), particularly parallel code pairs between Rust and other PLs, which are vital to code-to-code translation (Ahmad et al., 2023) while limited in the practical training corpus. In comparison, there may be more natural language-Rust code pairs, such as Rust documentation and programming tutorials, which help LLMs learn to generate Rust code matching natural language descriptions. The pseudocode-based translation enables LLMs to generate Rust code in this manner, which may better align with the training corpus. Also, translating from Rust (top row in heatmaps) gains noticeable improvement in many trials, suggesting that pseudocode may help the understanding of original programs written in Rust. The results demonstrate the helpfulness of pseudocode in assisting code translation involving a low-resource PL.
Message 3: Pseudocode-based translation generally improves code translation accuracy across most PL pairs. Besides, the improvement is more significant when translating from flexible and lightweight PLs to strict and complex PLs, and when translation involves the low-resource Rust. Practitioners are highly advised to consider pseudocode-based translation in these scenarios.
4.3. RQ3: Comparison with Programming Language as Transitive Intermediary
A recent study (Macedo et al., 2025) reveals that PLs themselves can also serve as effective intermediaries in transitive code translation to complement direct translation. For example, translating Python to Rust and then to Java helps resolve some translation tasks that cannot be handled by direct Python-to-Java translation (Macedo et al., 2025). Given PL as an effective baseline of transitive translation, this RQ verifies whether pseudocode can serve as a more general and effective transitive intermediary.
To conduct this comparison, we follow Macedo et al. (2025) to implement a baseline by substituting the transitive intermediary in our hybrid translation strategy (i.e., pseudocode) with a specific PL $L_{i}$, which is denoted as $\mathcal{D}\&\mathcal{PL}_{L_{i}}$. Specifically, given a source PL $L_{s}$, a target PL $L_{t}$, and an intermediate PL $L_{i}$ ($L_{i}\notin\{L_{s},L_{t}\}$), the baseline first translates the original program into an intermediate program in $L_{i}$, and then translates this intermediate program into $L_{t}$. The resulting program in $L_{t}$ is taken as the final translated result. The translations are performed by LLMs using the same prompt shown in Figure 3(a). Since no universally effective intermediate PL has been identified (Macedo et al., 2025), we implement the baseline for each studied PL and compare our $\mathcal{D}\&\mathcal{PC}$ strategy against each $\mathcal{D}\&\mathcal{PL}_{L_{i}}$, where $L_{i}\in\left\{\text{Python},\text{JavaScript},\text{C++},\text{Java},\text{Go},\text{Rust}\right\}$, with pass@10 as the metric.
(a) vs. $\mathcal{D}\&\mathcal{PL}_{\text{Python}}$ (Qwen-32B)
(b) vs. $\mathcal{D}\&\mathcal{PL}_{\text{JavaScript}}$ (Qwen-32B)
(c) vs. $\mathcal{D}\&\mathcal{PL}_{\text{C++}}$ (Qwen-32B)
(d) vs. $\mathcal{D}\&\mathcal{PL}_{\text{Java}}$ (Qwen-32B)
(e) vs. $\mathcal{D}\&\mathcal{PL}_{\text{Go}}$ (Qwen-32B)
(f) vs. $\mathcal{D}\&\mathcal{PL}_{\text{Rust}}$ (Qwen-32B)
Figure 5. Relative Improvement in Pass@10 Rates of Qwen-32B with Hybrid Strategy based on Pseudocode Compared to an Intermediate PL. (The greener, the more improvement. Red value means accuracy drop.)
Figure 5 presents the comparison results of pseudocode-based translation ($\mathcal{D}\&\mathcal{PC}$) versus intermediate PL-based translation ($\mathcal{D}\&\mathcal{PL}_{L_{i}}$) on Qwen32B. The results show that pseudocode-based translations achieve higher code translation accuracy than intermediate PL-based translations in most cases. Specifically, among all 120 comparisons ($5$ source PLs $\times$ 44 target PLs $\times$ 66 transitive PLs), pseudocode-based translation wins in 109 comparisons (green cells), whereas it loses in only 11 comparisons (red cells) with marginal differences (less than 1%). Although certain intermediate PLs are more helpful than pseudocode for some source-target PL pairs (e.g., using JavaScript as the intermediate PL for Rust-to-Python, Go-to-C++, Go-to-Java, and C++-to-Java translation, Figure 5(b)), the pseudocode-based translations yield a generally more pronounced outperformance. The observed trends and conclusions are consistent across the other experimental LLMs (i.e., Qwen7B, Phi-4, GPT-4o-mini, and DSCoder). To avoid redundancy, we omit their detailed results here. These results are available in our artifact (Artifact of this paper, [n. d.]).
Through case studies, we found that the advantage of pseudocode over PLs is its role as a general abstraction to bridge differences across diverse source-target PL pairs. Specifically, aligning with findings in (Macedo et al., 2025), we observed that a transitive PL is effective when it bridges the different features and APIs between the source and target PLs; meanwhile, different PL pairs may require varying transitive PLs. An inappropriate choice of intermediate PL may even increase the gap between the source and target PLs and introduce extra burdens in code translation, hindering the translation accuracy. Besides, an intermediate PL may also bring LLMs new burdens to satisfy its syntax and semantics, complicating the translation process. In contrast, flexible and descriptive pseudocode describes code intent and logic, offering a PL-agnostic abstraction that guides LLMs to focus on semantics instead of detailed PL-specific implementations, thereby offering better effectiveness in general. We also examined cases where pseudocode underperformed PLs in transitive translation and found that inferiority mainly stemmed from the missing essential details in the pseudocode (which will be discussed in Section 4.5, Limit-2). In comparison, programs in intermediate PLs often preserve the complete code semantics, facilitating the generation of a semantically consistent translation result. These findings highlight the advantages of pseudocode as a general effective intermediary, as well as the need for enhancing its quality to fully harness its effectiveness.
Message 4: Pseudocode serves as a more general and effective intermediary than a specific PL in transitive code translation for most studied PL pairs. This is because pseudocode provides a PL-agnostic abstraction of code semantics that generally bridges differences across diverse source-target PL pairs, as an effective transitive PL acts for their fitting PL pairs.
4.4. RQ4: Effectiveness of Code Translation based on High-Quality Pseudocode
The pseudocode generated by the studied LLMs may introduce noise (e.g., incorrect or ambiguous descriptions of code intent and logic, which will be discussed in Section 4.5). Such low-quality pseudocode can mislead code translation. This RQ therefore examines whether better performance can be achieved when high-quality pseudocode is provided and identifies the bottlenecks in pseudocode-based code translation. Specifically, Wu et al. (2025) demonstrate that DeepSeek-R1 can help annotate human-written-like pseudocode with high accuracy and naturalness for concrete programs. Thus, in this RQ, we take DeepSeek-R1 as a high-quality pseudocode generator [^1] To avoid potential data contamination of DeepSeek-R1, whose knowledge cutoff date is unknown, we only adopt it to prepare high-quality pseudocode following Wu et al. (2025). We do not include it as a translator LLM to study.. Then, we evaluate the translation accuracy of the studied LLMs based on DeepSeek-R1-generated high-quality pseudocode. The investigation results help indicate the potential effectiveness of pseudocode-based translation when higher-quality pseudocode is available (e.g., written by humans or generated by enhanced pseudocode generation methods).
Table 3. Pass@10 of Code Translation Based on Pseudocode Generated by Translator LLM (Self-Gen) and DeepSeek-R1 (R1-Generated).
LLM Strategy Easy Tasks Medium Tasks Hard Tasks
Self-Gen R1-Generated Self-Gen R1-Generated Self-Gen R1-Generated
Qwen32B $\mathcal{D}$ 0.9480 - - 0.8772 - - 0.7059 - -
$\mathcal{P}$ 0.9442 0.9569 (+1.35%) 0.8736 0.8971 (+2.69%) 0.6768 0.7376 (+8.98%)
$\mathcal{PC}$ 0.9648 0.9718 (+0.73%) 0.9001 0.9227 (+2.51%) 0.7553 0.7858 (+4.04%)
$\mathcal{D}\&\mathcal{P}$ 0.9773 0.9833 (+0.61%) 0.9403 0.9512 (+1.16%) 0.7948 0.8260 (+3.93%)
$\mathcal{D}\&\mathcal{PC}$ 0.9773 0.9835 (+0.63%) 0.9327 0.9476 (+1.60%) 0.7988 0.8286 (+3.73%)
Phi-4 $\mathcal{D}$ 0.9276 - - 0.8313 - - 0.6281 - -
$\mathcal{P}$ 0.8782 0.9287 (+5.75%) 0.7744 0.8402 (+8.50%) 0.5832 0.6433 (+10.31%)
$\mathcal{PC}$ 0.9254 0.9572 (+3.44%) 0.8347 0.8796 (+5.38%) 0.6463 0.6986 (+8.09%)
$\mathcal{D}\&\mathcal{P}$ 0.9646 0.9781 (+1.40%) 0.8947 0.9192 (+2.74%) 0.7083 0.7563 (+6.78%)
$\mathcal{D}\&\mathcal{PC}$ 0.9610 0.9759 (+1.55%) 0.8903 0.9127 (+2.52%) 0.7038 0.7385 (+4.93%)
GPT-4o -mini $\mathcal{D}$ 0.9279 - - 0.8552 - - 0.6667 - -
$\mathcal{P}$ 0.8647 0.9448 (+9.26%) 0.7219 0.8538 (+18.27%) 0.5000 0.6752 (+35.04%)
$\mathcal{PC}$ 0.9239 0.9641 (+4.35%) 0.8635 0.8973 (+3.91%) 0.6837 0.7333 (+7.25%)
$\mathcal{D}\&\mathcal{P}$ 0.9658 0.9787 (+1.34%) 0.9054 0.9338 (+3.14%) 0.7404 0.7972 (+7.67%)
$\mathcal{D}\&\mathcal{PC}$ 0.9621 0.9790 (+1.76%) 0.9144 0.9327 (+2.00%) 0.7617 0.7915 (+3.91%)
Qwen7B $\mathcal{D}$ 0.9073 - - 0.7928 - - 0.5400 - -
$\mathcal{P}$ 0.7735 0.9216 (+19.15%) 0.5928 0.8042 (+35.66%) 0.3286 0.5390 (+64.03%)
$\mathcal{PC}$ 0.8989 0.9497 (+5.65%) 0.7929 0.8575 (+8.15%) 0.5480 0.6426 (+17.26%)
$\mathcal{D}\&\mathcal{P}$ 0.9336 0.9646 (+3.32%) 0.8395 0.8831 (+5.19%) 0.5849 0.6598 (+12.81%)
$\mathcal{D}\&\mathcal{PC}$ 0.9420 0.9638 (+2.31%) 0.8568 0.8861 (+3.42%) 0.6154 0.6747 (+9.64%)
DSCoder $\mathcal{D}$ 0.8964 - - 0.7876 - - 0.5586 - -
$\mathcal{P}$ 0.8562 0.9170 (+7.10%) 0.7335 0.7979 (+8.78%) 0.5069 0.5752 (+13.47%)
$\mathcal{PC}$ 0.9218 0.9509 (+3.16%) 0.7875 0.8540 (+8.44%) 0.5530 0.6504 (+17.61%)
$\mathcal{D}\&\mathcal{P}$ 0.9482 0.9678 (+2.07%) 0.8569 0.8916 (+4.05%) 0.6442 0.6934 (+7.64%)
$\mathcal{D}\&\mathcal{PC}$ 0.9537 0.9726 (+1.98%) 0.8581 0.8954 (+4.35%) 0.6459 0.7073 (+9.51%)
Average $\mathcal{D}$ 0.9214 - - 0.8288 - - 0.6199 - -
$\mathcal{P}$ 0.8634 0.9338 (+8.16%) 0.7392 0.8386 (+13.45%) 0.5191 0.6341 (+22.15%)
$\mathcal{PC}$ 0.9270 0.9587 (+3.43%) 0.8357 0.8822 (+5.56%) 0.6373 0.7021 (+10.18%)
$\mathcal{D}\&\mathcal{P}$ 0.9579 0.9745 (+1.73%) 0.8874 0.9158 (+3.20%) 0.6945 0.7465 (+7.49%)
$\mathcal{D}\&\mathcal{PC}$ 0.9592 0.9750 (+1.64%) 0.8905 0.9149 (+2.74%) 0.7051 0.7481 (+6.10%)
(+xx%): performance improvement of a strategy based on R1-generated pseucode compared to LLM-translator-generated pseudocode.
Improvement with High-quality Pseudocode. Table 3 compares the pass@10 rates of different translation strategies based on high-quality pseudocode generated by DeepSeek-R1 and pseudocode generated by the studied LLMs themselves (i.e., the default setup studied in previous RQs), respectively. The results show that the code translation accuracy consistently improves across LLMs and strategies when using R1-generated pseudocode. Among all strategies, the pure pseudocode-based strategy ($\mathcal{P}$) benefits the most, with average improvements of 8.16%, 13.45%, and 22.15% on three difficulty levels, respectively, mitigating the underperformance relative to direct translation ($\mathcal{D}$) observed in RQ1. These results indicate pseudocode quality as a critical bottleneck for pseudocode-based translation. In addition, we observed that hybrid strategies ($\mathcal{D}\&\mathcal{P}$ and $\mathcal{D}\&\mathcal{PC}$) continuously outperform single-approach strategies ($\mathcal{D}$, $\mathcal{P}$, $\mathcal{PC}$), indicating that combining direct and pseudocode-based translation remains beneficial even with higher-quality pseudocode. With R1-generated high-quality pseudocode, the best pass@10 scores of all LLMs with optimal strategy improved to 0.9646-0.9835, 0.8861–0.9512, and 0.6747–0.8286 across the three difficulty levels, demonstrating the promising potential of pseudocode-based code translation.
Message 5: The quality of pseudocode hinders accurate pseudocode-based translation of the studied LLMs. Higher-quality pseudocode can consistently improve the code translation accuracy of studied LLMs across translation strategies and task difficulties.
LLM-wise Comparison. The performance comparison among LLMs further demonstrates the remaining limitations of LLMs in both code understanding and generation during pseudocode-based code translation. Specifically, we observed that the performance improvement brought by R1-generated pseudocode is more significant for weaker LLMs. For example, the improvement of all strategies on Qwen7B ranges from 2.31% to 64.03%, while it is only 0.61% to 8.98% on Qwen32B, indicating that higher-quality pseudocode is more beneficial for weaker LLMs. This demonstrates the weakness of weaker LLMs in understanding and summarizing the code intent and logic of original programs. In addition, with R1-generated pseudocode, the performance gap between weaker LLMs and more powerful LLMs remains. For example, with R1-generated pseudocode, the pass@10 rates of Qwen7B and DSCoder with $\mathcal{D}\&\mathcal{PC}$ are only 0.6747 and 0.7073 on hard tasks, respectively, which are much lower than 0.8286 of Qwen32B. The results indicate that even given the same code intent and logic (i.e., pseudocode), the weaker LLMs are still less capable in code translation than the more powerful LLMs, indicating that the code implementation capability also hinders the effectiveness of pseudocode-based code translation on weaker LLMs.
Message 6: Both the code understanding capability (indicated by the quality of the generated pseudocode) and the code implementation capability of the studied LLMs hinder their performance in code translations.
4.5. Discussion: Advantages and Limitations of Pseudocode-Based Translation
After understanding the effectiveness of pseudocode-based code translation from the quantitative results in four RQs, we further conduct case studies to attribute its successes and failures to learn concrete insights for the effective adoption of this approach. Based on these findings, we also suggest several research directions to mitigate the limitations and harness the potential.
Key advantages of pseudocode-based code translation. We identified three major advantages of emulating explicit semantic translation via pseudocode-based translation to complement direct translation. One of them facilitates the understanding of the original programs and reduces the burden of one step, and the other two help LLMs handle differences across PLs. Practitioners are encouraged to leverage the advantages of pseudocode-based translation for original programs or PL pair fitting them.
  1. Pro-1
    Pseudocode helps disentangle code understanding and implementation in code translation into two steps, formulating easier subproblems for LLMs to solve over complicated programs.
Some programs are compact, implementing multiple operations in a single dense statement by chaining several steps recursively or sequentially (Keshk and Dyer, 2023). Such programs may result from PL idioms or developers’ programming style (Keshk and Dyer, 2023). We observed that LLMs struggle to disentangle the functionality of such complicated programs and re-implement the functionality in the target PL via one-step direct translation. LLMs often miss or mess up steps in such cases. In comparison, a pseudocode-based transitive translation splits the code understanding of the complicated program and the implementation of the code intent and logic into two steps, reducing the burden on LLMs and promoting accuracy in each step.
Figure 6 illustrates an example of this case. Specifically, the original program includes only one Python statement that recursively conducts four operations (LABEL:lst:_rq1-type3-py2js-python). When translating this compact program via the one-step direct translation, even the powerful Qwen32B missed the row-sorting step in the output JavaScript program (LABEL:lst:_rq1-type3-py2js-direct). In comparison, when asked to describe the code intent and logic using pseudocode, Qwen32B was able to interpret the complicated logic in the compact Python program in pseudocode (LABEL:lst:_rq1-type3-py2js-pseudocode). The pseudocode then promotes the successful implementation of the operation JavaScript (LABEL:lst:_rq1-type3-py2js-viapseudocode). Similar symptoms are also observed on Phi-4 and Qwen7B.
def matrixSum(self, nums: List[List[int]]) -> int:
return sum(max(col) for col in zip(*[sorted(row) for row in nums]))
# organizing row-wise sorting, transposition, max
# val extraction, and sum in a single expression.
Listing 1: Source Python Program to Translate
function matrixSum(nums)
sorted_rows = sort each row in nums
transposed_matrix = transpose sorted_rows
max_values = find the maximum value in each col...
return sum of max_values
Listing 2: Pseudocode of Source Python Program
var matrixSum = function(nums) {
return nums[0].map((_, colIndex) => Math.max(
...nums.map(row => row[colIndex]))//unsorted rows
).reduce((sum, maxVal) => sum + maxVal, 0);
Listing 3: Translation with Direct
var matrixSum = function(nums) {
const sortedRows = nums.map(row => row.sort((a, b) => a - b)); // sort row before other ops
const transposedMatrix = sortedRows[0].map(...
Listing 4: Translation with viaPseudocode (w/ code)
Figure 6. An Example of Python-to-JavaScript Translation Involving Compact Original Programs (Qwen32B)3
  1. Pro-2
    Pseudocode abstracts the PL-specific features (e.g., variable scoping, closure capture mode, and syntax sugar), avoids LLMs in following the incompatible practices in original programs.
We observed that LLM-based direct translation may overlook the differences in the features of source and target PLs, resulting in incompatible implementations violating the mechanism of the target PL. Meanwhile, pseudocode describes the code’s intent and logic, hiding PL-specific implementation. When generating target code from the pseudocode, LLMs often properly implement the intended function in a manner fitting the mechanism of the target PL, and are less likely to be distracted by the incompatible patterns of source PL in the original program.
Footnote 2 shows an example of this case with a Go-to-Rust translation. Specifically, when translating a Go program with an inner function (LABEL:lst:_rq1-type1-go2rust-go) into Rust via direct translation, Qwen7B mechanically followed the Go program to declare a common inner function in Rust without realizing the difference in default closure capture modes in Go and Rust and forgetting to indicate a mutable closure capture mode (LABEL:lst:_rq1-type1-go2rust-direct). Meanwhile, based on pseudocode that mainly indicates the code intent and logic (LABEL:lst:_rq1-type1-go2rust-pseudocode), Qwen7B correctly implemented the functionality in a proper manner that fits Rust’s memory management (LABEL:lst:_rq1-type1-go2rust-viapseudocode). Interestingly, all five studied LLMs failed with the same fault in direct translation while succeeding with the help of pseudocode. This also echoes the helpfulness of pseudocode-based code translation in translating from flexible PLs to more rigid PLs observed in RQ2 (Section 4.2).
func canMakeSquare(grid [][]byte) bool {
var res int
calcRes := func (r byte) { // mutable closure
if r == ’W’ { res++; } //capture by default
Listing 5: Source Go Program to Translate
function canMakeSquare(grid):
count = count W in the 2x2 block
Listing 6: Pseudocode of Source Go Program
pub fn can_make_square(grid: Vec<Vec<char>>) -> bool {
let mut res = 0;
let calc_res = |r: char| { //mimic Go code, but
if r == ’W’ {//by default immutable closure
res += 1;//, thus illegal to mutate res
Listing 7: Translation with Direct
pub fn can_make_square(grid: Vec<Vec<char>>) -> bool {
let mut count = 0;
// flexibly count without involving closure
for x in 0..2 { for y in 0..2 {
if grid[i + x][j + y] == ’W’ {
count += 1;
Listing 8: Translation with viaPseudocode (w/ code)
Figure 7. An Example of Go-to-Rust Translation Involving Unique Closure Capturing in Go and Rust (Qwen7B)333“…” represents code less related to the illustration and omitted for saving space. The comment is not generated by LLMs but annotated by authors for case explanation.
  1. Pro-3
    Pseudocode describes the functionality implemented by unique APIs in source PL, allowing LLMs to generate flexible implementations to realize the intended functionality in target PL.
We observed that LLMs could hallucinate incorrect implementations that fail to realize the functionality of the PL-specific APIs in original programs during direct translation. LLMs struggle to identify or implement a semantically equivalent API counterpart in the target PL based on only the original program, and may even hallucinate non-existing APIs. Meanwhile, pseudocode often describes the functionality implemented with PL-specific APIs in natural language, guiding LLMs to figure out appropriate APIs and API usages in the target PL as they work out a code generation task without being misled by concrete API usages in original programs.
Figure 8 illustrates an example of this case. Specifically, the original Python program uses an API “bit_count” to calculate the number of ones in the binary expression of an integer (LABEL:lst:_rq1-type2-py2go-python). Due to no “bit_count” API in Go, when translating the Python program into Go via direct translation, Phi-4 implemented bit counting with a series of bit operations (LABEL:lst:_rq1-type2-py2go-direct). However, the generated bit operations do not align with the functionality of “bit_count”. Similarly, GPT-4o-mini hallucinated incorrect bit operations, and DSCoder even hallucinated a non-existing API. Instead, the pseudocode generated by Phi-4 clarified the functionality implemented by “bit_count” in Python (LABEL:lst:_rq1-type2-py2go-pseudocode), which guided Phi-4 to generate a correct subroutine “countBits” in Go to realize the functionality of “bit_count” (LABEL:lst:_rq1-type2-py2go-viapseudocode). DSCoder and GPT-4o-mini also generated a correct translation based on the pseudocode generated by themselves.
def kthCharacter(self, k: int) -> str:
return ascii_lowercase[(k-1).bit_count()]
Listing 9: Source Python Program to Translate
function kthCharacter(k)
index = count the number of 1s in the binary representation of (k-1)
return character at index from string_of_lowercase_letters
Listing 10: Pseudocode of Source Python Program
func kthCharacter(k int) byte {
return ascii_lowercase[(k-1)^((k-1)&-((k-1)))-1]
// Calculating a mask instead of counting 1 in the binary expression.
// We leverage the space below to show hallucinated implementation of GPT-4o-mini and DSCoder.
// GPT-4o-mini:
// return ascii_lowercase[(k-1)&0x1F] //diff func.
// DSCoder:
// index := (k - 1).bit_count() //non-existing API
Listing 11: Translation with Direct
func kthCharacter(k int) byte {
index := countBits(k - 1)
return stringOfLowercaseLetters[index]
func countBits(n int) int {
count := 0
for n > 0 {
count += n & 1
n >>= 1
return count
Listing 12: Translation with viaPseudocode (w/ code)
Figure 8. An Example of Python-to-Go Translation Involving Non-Existing Equivalent APIs in Go (Phi-4)3
Typical limitations of pseudocode-based code translation. We also observed three major limitations of pseudocode that cause errors in semantic code translation results. They stem from LLMs’ incorrect understanding of original programs, the semantic loss in information transmission, and the ambiguity of natural language. Enhancements to generate more accurate pseudocode are needed to mitigate these issues and harness the advantages of pseudocode-based code translation.
  1. Limit-1
    Pseudocode generated by LLMs may be incorrect, describing an intent or logic inconsistent with those of the original program and misleading the implementation in the target PL.
Explicit semantic code translation based on pseudocode introduces a compulsory code understanding step during translation. However, LLMs are not free from erroneous code understanding. An incorrect code intent or logic can conversely disturb code translation. For example, translating a straightforward Python expression “sum = sum+nums[i] if nums[i]<=sum else nums[i]” into Java may not necessitate the understanding of the code functionality. An incorrect intent conversely misled translation: Qwen32B hallucinated a code intent of “sum = max(sum, sum + nums[i])” for this expression, misleading the generation of the Java program. The issue happened with both $\mathcal{P}$ and $\mathcal{PC}$ strategies, while direct translation ($\mathcal{D}$) leads to a correct translation.
  1. Limit-2
    Pseudocode generated by LLMs may miss information essential to the reproduction of the complete functionality of the original program.
LLM-generated pseudocode may also miss certain essential information for implementing the complete functionality of the original program. For example, the data type of variables is often omitted in the LLM-generated pseudocode without any description of precision concerns, leaving operations sensitive to data type (e.g., requiring certain precision) hard to reproduce. Another typical missing information for reproducing code semantics is the logic of complicated algorithms. From a concise description without detailed implementations, LLMs may not accurately reproduce all algorithms. The concise description may also miss customized operations in algorithms, leaving the generated code degenerate into a general implementation lacking the expected features.
  1. Limit-3
    Natural language description in pseudocode may induce ambiguity in elaboration of code intent and logic.
Although natural language can depict PL-agnostic code intent and logic, its nature of ambiguity may introduce noise and mislead the code generation step. A typical issue observed in the failure cases is the ambiguity of “to” and “downto” in “for” loops. Such expressions cannot clarify the inclusion of the border value in the loop. The inclusion of the original program (i.e., using $\mathcal{PC}$ strategy) cannot always mitigate the disturbance caused by the ambiguity as well. In comparison, a direct translation from the original program does not suffer from such ambiguity-induced issues.
Research Opportunities. To leverage the advantages of pseudocode-based code translation observed in the previous three RQs and its potential revealed in RQ4, we suggest four directions for future enhancement based on our findings. These target the generation and validation of pseudocode, as well as hybrid approaches to combine the strengths of direct and transitive translation.
  1. Chance-1
    Refining pseudocode generation through more fine-grained rules that are tailored for code translation tasks.
In this study, we rely on generic instructions instructed by Wu et al. (2025). Incorporating specific guidelines may help guide LLMs to preserve essential information and minimize ambiguities. The refinement may help capture essential precision requirements, algorithmic details, and clear control logic like loop boundaries that are frequently omitted in standard pseudocode generation, mitigating the information loss and ambiguity that we observed in Limit-2 and Limit-3.
  1. Chance-2
    Designing systematic validation and repair mechanisms for generated pseudocode.
Effective validation strategies may help identify inconsistencies and semantic losses. For example, round-trip validation (Aiken and Park, 2010; Zhou and Sun, 2018) may identify buggy pseudocode by translating the generation result back to the source PL, and LLMs may infer the intended behavior based on code (Xu et al., 2025) and compare pseudocode to it to debug pseudocode. Incorporating such validation loops would help identify and mitigate failures stemming from low-quality pseudocode suffering from the three limitations.
  1. Chance-3
    Automating selection of the translation strategy for each task without dynamic validation.
In this study, we collect the generation results of both strategies and rely on dynamic test execution to determine the optimal strategy. A lightweight static classifier or heuristic that inspects source code characteristics and pseudocode quality for early selection or rejection of translation strategies could streamline the translation process and reduce computational overhead. In addition, the advancement in LLM-based execution prediction may also facilitate validation of candidate translations without dynamic execution (Xu et al., 2025).
  1. Chance-4
    Combining pseudocode-based and direct translations with mutual information.
Exploring more seamless combination approaches of the two approaches may also further enhance the resulting code translation accuracy. For example, pseudocode could be leveraged to guide the validation and repair of direct translation results, or vice versa. Such integration may help integrate the fine-grained advantages of both approaches.
5. Threats to Validity
Table 4. Pass@10 Rates on All 323 Tasks and 114 Newer Tasks After Cut-off Date (2024-06-01)
Strategy LLM All 323 Tasks
114 Newer Tasks
LLM All 323 Tasks
114 Newer Tasks
$\mathcal{D}$(Direct Translation) Qwen 32B 0.8777 0.8604 Qwen 7B 0.7934 0.7623
$\mathcal{P}$(viaPseudocode) 0.8703 (-0.84%) 0.8424 (-2.09%) 0.7446 (-21.43%) 0.5320 (-30.21%)
$\mathcal{PC}$(viaPseudocode w/ prog.) 0.9023 (+2.80%) 0.8848 (+2.84%) 0.8016 (+2.96%) 0.7435 (-2.47%)
$\mathcal{D}\&\mathcal{P}$(Hybrid [DT&vP]) 0.9325 (+6.24%) 0.9199 (+6.92%) 0.8587 (+4.92%) 0.7956 (+4.37%)
$\mathcal{D}\&\mathcal{PC}$(Hybrid [DT&vPw/p]) 0.9292 (+5.87%) 0.9170 (+6.58%) 0.8616 (+6.93%) 0.8151 (+6.93%)
$\mathcal{D}$(Direct Translation) Phi-4 0.8363 0.8146 DS Coder 0.7934 0.7672
$\mathcal{P}$(viaPseudocode) 0.7839 (-6.27%) 0.7702 (-5.45%) 0.7446 (-5.23%) 0.7185 (-6.35%)
$\mathcal{PC}$(viaPseudocode w/ prog.) 0.8398 (+0.42%) 0.8210 (+0.79%) 0.8016 (+3.30%) 0.7792 (+1.56%)
$\mathcal{D}\&\mathcal{P}$(Hybrid [DT&vP]) 0.8926 (+6.73%) 0.8773 (+7.70%) 0.8587 (+8.23%) 0.8330 (+8.58%)
$\mathcal{D}\&\mathcal{PC}$(Hybrid [DT&vPw/p]) 0.8886 (+6.25%) 0.8735 (+7.23%) 0.8616 (+8.60%) 0.8359 (+8.95%)
$\mathcal{D}$(Direct Translation) GPT-4o -mini 0.8539 0.8363 Avg 0.8309 0.8082
$\mathcal{P}$(viaPseudocode) 0.7409 (-13.23%) 0.7155 (-14.44%) 0.7769 (-9.04%) 0.7157 (-11.44%)
$\mathcal{PC}$(viaPseudocode w/ prog.) 0.8590 (+0.60%) 0.8520 (+1.88%) 0.8409 (+2.48%) 0.8161 (+0.98%)
$\mathcal{D}\&\mathcal{P}$(Hybrid [DT&vP]) 0.9031 (+5.76%) 0.8857 (+5.91%) 0.8891 (+6.37%) 0.8623 (+6.70%)
$\mathcal{D}\&\mathcal{PC}$(Hybrid [DT&vPw/p]) 0.9093 (+6.49%) 0.8962 (+7.16%) 0.8901 (+6.80%) 0.8675 (+7.35%)
We discuss three potential threats to the validity of our study and our mitigation methods as follows.
Representativeness of Studied Subjects. The first threat to our study is about the representativeness and generalizability of the observations on our study subjects. To mitigate this threat, we followed existing work to investigate code translation across six PLs. These include Python, C++, Java, and JavaScript, which are widely adopted in daily development and studied in existing works (TIOBE Index, [n. d.]; Rozière et al., 2022; Jiao et al., 2023; Yang et al., 2024), as well as the emerging Go and Rust (Pan et al., 2024; Macedo et al., 2025). Our study results are expected to guide the adoption of pseudocode-based translation for these popular PLs.
We constructed translation tasks based on LeetCode problems in LiveCodeBench (Jain et al., 2025). LeetCode problems are widely used in code translation studies (Wang et al., 2023a; Ahmad et al., 2023) and benchmark of LLMs’ coding ability (Hui et al., 2024; Team, 2024; Jain et al., 2025) since they provide adequate test cases to evaluate candidate solutions and have easy-to-collect multilingual solution programs. In addition, these tasks require common engineering and algorithm implementations useful in daily development. Thus, we consider the study results on their solution programs should be meaningful and generalizable to practical development code.
We used five LLMs deployable on our machine as the code translator. These include both general LLMs and coder LLMs, and both closed-source and open-source LLMs from four families. They rank a varying range (from Top-10 to 68) on the famous BigCodeBench (Zhuo et al., 2025). Also, the training data for these LLMs has a cutoff date with relatively little overlap with the timeline of LeetCode problems. Thus, we consider the results learned with these LLMs should be representable and reliable.
Data Contamination. A potential threat to our study result is the subject to the data contamination issue, where LLMs perform well because they have learned the ground truth during training rather than pseudocode-based translation strategies. Although our translation tasks based on LiveCodeBench are much newer than the subjects in the conventional code translation benchmarks as introduced in Section 3.2, there is still some overlap between the timeline of LiveCodeBench problems and the studied LLMs’ training data. To further verify the cause of the observed performance improvement, we re-evaluate LLMs on the solution programs of a clean subset of 114 LeetCode problems released after the latest knowledge cutoff of the studied LLMs, i.e., June 1, 2024. As shown in Table 4, the performance improvement of different strategies is generally consistent with that observed on all 323 tasks. Thus, we consider the observation based on the complete set is meaningful and provides more statistically meaningful observations with more subjects.
Representativeness of Studied Translation Strategies. The representativeness of the studied translation strategies and prompts is another potential threat to our study, impacting the meaningfulness of our observations. To mitigate this threat, we studied five translation strategies, including one direct translation strategy, which is the commonly used approach in existing studies and practices, as well as two pseudocode-based translation strategies and two hybrid translation strategies. The pseudocode-based strategies emulate the human practices of semantic translation on natural languages (Newmark, 1981) and are implemented in a transitive translation manner following (Macedo et al., 2025). The hybrid strategies are based on straightforward result combination to explore the helpfulness of the combined advantages of strategies, considering the feasibility of selecting multiple translation candidates based on original programs (Macedo et al., 2025). We carefully designed the prompts following existing studies of direct code translation (Yang et al., 2024; Pan et al., 2024) and pseudocode-based code generation (Wu et al., 2025). Thus, they can reflect the typical usage of these strategies and prompts and lead to meaningful observations.
6. Related Work
6.1. Automated Code Translation
To facilitate development activities relying on code translation, various automated code translation methods have emerged in the past decades (Chen et al., 2025; Jiao et al., 2023). Nguyen et al. (2013) pioneered in automated code translation via statistical machine translation. To improve translation accuracy by operating directly on ASTs or program graphs, researchers designed tree- and graph-based neural models (Chen et al., 2018; Guo et al., 2021). Unsupervised and self-supervised model training approaches were later proposed to mitigate the scarcity of parallel code corpora (Rozière et al., 2020, 2022; Liu et al., 2023).
Given the impressive capabilities of LLMs in various coding tasks (Zhang et al., 2023), recent works have explored LLMs for code translation. Yang et al. (2024) demonstrated that LLMs can serve as effective code translators. Pan et al. (2024) summarized the common errors made by LLMs in code translation, aiming to inspire future enhancement of LLM-based code translation. Ibrahimzada et al. (2025) leveraged LLMs for project-level code translation. These works mainly focused on direct translation from source PL to target PL, where LLMs are fed with the original program and output the translated program. Recently, Macedo et al. (2025) found that a one-step direct translation from source PL to target PL may not leverage the full potential of LLMs due to the varying gap between source and target PLs. They proposed transitive direct translations with an intermediary PL as a bridge.
Existing automated code translation approaches mainly focus on code-to-code transformation. Differently, we study the effectiveness of emulating semantic translation for code via pseudocode, where LLMs first describe the semantics of original programs with pseudocode and then produce translated programs. We identify the complementary role of pseudocode-based translation for conventional direct translation, as well as its helpfulness in reducing task difficulties and bridging the differences between the features and APIs in varying PL pairs. The findings may guide practitioners to obtain accurate LLM-generated code translations by leveraging pseudocode.
In addition, some works attempted to debug the erroneous code translations, e.g., by locating code snippets leading to inconsistent behaviors between the original and translated programs to guide repair (Wang et al., 2023b) and conducting certain types of repair based on fixed templates (Ramos et al., 2024). The methods mainly debug minor mistakes in the translated code. They are generally orthogonal to our pseudocode-based code translation approach, which aims to reduce the occurrence of errors in the translation results and may also facilitate patch generation. Our methods can be combined with the debugging methods to further enhance the accuracy of LLM-generated code translations.
6.2. Applications of Pseudocode
Existing works exploit pseudocode for diverse code-related tasks. A popular task is to generate concrete programs from pseudocode. Dirgahayu et al. (2017) designed an XML-based method to help novice programmers learn PLs based on programs generated from pseudocode. Kulal et al. (2019) proposed a search-based method and Acharjee et al. (2022) applied Seq2Seq neural models to enhance the code generation performance. Pseudocode is also used as an intermediate representation to enhance multilingual code generation accuracy (Sun et al., 2024). Recently, Wu et al. (2025) annotated pseudocode for LeetCode problems to isolate the evaluation of problem-solving and language-coding abilities of LLMs in code generation. They also reveal that LLMs can generate programs in different PLs based on pseudocode and synthesize pseudocode for concrete programs, which motivates our study.
In addition to code generation, pseudocode has also proven useful in cross-PL code retrieval by embedding structured pseudocode representations (Kulkarni et al., [n. d.]), in binary function similarity and vulnerability search by leveraging decompiled pseudocode to obtain platform-agnostic semantic representations (Zhang et al., 2022), and in generating descriptive summaries for stripped binaries using pseudocode extracted by decompilers as an expert-guided signal (Ye et al., 2023). All these applications leveraged the advantage of pseudocode in succinctly capturing the intent and logic of programs to facilitate code comprehension and semantic representation (Oda et al., 2015).
This work studies the helpfulness of pseudocode for another common software development task, i.e., code translation, where pseudocode serves as a semantic interpretation to facilitate semantic translation. The extensive investigation of pseudocode’s effectiveness in code translation yields various findings, which may guide practitioners to obtain accurate LLM-generated code translations based on pseudocode. We also identified limitations of pseudocode generated by LLMs, which are expected to shed light on the focus of validation and repair for automatically generated pseudocode, helping harness the potential of pseudocode for downstream tasks like code translation.
7. Conclusion
In this work, we empirically study the effectiveness of pseudocode-based code translation with LLMs, which explicitly emulates the human practice of semantic translation. By investigating the performance of five popular LLMs with pseudocode-based code translation on 9,690 translation tasks across six popular PLs, we reveal that pseudocode-based code translation can effectively complement the widely adopted direct code translation approach for various pairs of source and target PLs. We also demonstrate the effectiveness of pseudocode as a general and effective intermediary, as well as the further potential of translation based on higher-quality pseudocode and the bottleneck in both code understanding and generation. Our case studies further identify the concrete advantages and limitations of pseudocode-based code translation. Based on these findings, we suggest the hybrid use of pseudocode-based and direct code translation to enhance code translation accuracy, as well as discuss future research directions to further unleash the potential of pseudocode in code translation.

### v3

@@
 Abstract.
-Large language models (LLMs) show great potential in code translation.
-However, accurate translation remains challenging when using the commonly adopted direct code-to-code translation approach, which converts a program into the target programming language (PL) in a single step.
-Inspired by the success of incorporating intermediate steps to guide LLMs in resolving challenging tasks, we explore pseudocode-based code translation, which emulates the human semantic translation by first interpreting the program’s intent and logic into pseudocode and then implementing it in the target PL.
-We find that pseudocode-based translation helps translate programs that direct translation struggles to handle.
-Nonetheless, the effectiveness, advantages, and limitations of this approach remain underexplored.
-To bridge this gap, we present an empirical study on pseudocode-based code translation, aiming to investigate its effectiveness in enhancing the direct translation approach, illuminate its effective usage, and identify limitations hindering its potential benefits.
-By comparing direct and pseudocode-based translation approaches on 9,690 translation tasks across six PLs with five popular LLMs, we demonstrate that pseudocode-based translation can effectively complement direct translation, particularly when translating from flexible to rigid PLs or dealing with low-resource Rust.
-Based on these findings, we suggest adopting strategies that combine the complementary strengths of both approaches to enhance code translation accuracy.
-We also reveal the advantages of pseudocode-based translation in disentangling translations of complicated programs and mitigating distractions from detailed implementations in original programs, as well as its limitations due to incorrect, incomplete, or ambiguous pseudocode.
+Although large language models (LLMs) show promising potential in code translation, they still struggle to generate accurate translations using the commonly adopted direct code-to-code translation approach, which converts an original program into the target programming language (PL) in a single step.
+Inspired by the success of incorporating intermediate steps to guide LLMs in resolving challenging tasks, in this study, we explore pseudocode-based code translation.
+This approach emulates human semantic translation by first interpreting the original program’s intent and logic into pseudocode and then implementing it in the target PL.
+To understand the effectiveness of this underexplored approach, we present a systematic empirical study on pseudocode-based code translation, aiming to investigate its helpfulness in enhancing the direct translation approach, illuminate its effective usage, and identify its limitations.
+By comparing direct and pseudocode-based translation on 9,690 translation tasks across six PLs with five popular LLMs, we found that pseudocode-based translation can effectively complement direct translation, particularly when translating from flexible to rigid PLs and handling a low-training-resource PL.
+Based on the findings, we suggest combining the translation results of both approaches for test-based selection to leverage their complementary strengths.
+We also reveal the advantages of pseudocode-based translation in decoupling the code understanding and generation burden on complicated programs and mitigating distractions from PL-specific implementations in original programs, as well as its limitations due to incorrect, incomplete, or ambiguous pseudocode.
+Our study sheds light on the effective use of pseudocode-based translation and provides evidence to help enhance LLMs in code translation.
 Code Translation, Pseudocode, Semantic Translation, Large Language Model
+copyright: ccjournal: TOSEMjournalyear: 2026journalvolume: 1journalnumber: 1publicationmonth: 1doi: 10.1145/3790101ccs: General and reference Empirical studiesccs: Software and its engineering Software maintenance tools
 1. Introduction
-Code translation, also known as transpilation, refers to the automatic conversion of a program written in one programming language (PL) to another while preserving its semantics (i.e., program functionality) (Nguyen et al., 2013; Rozière et al., 2020; Pan et al., 2024).
+Code translation, also known as transpilation, refers to the automatic conversion of a program written in one programming language (PL) to another while preserving its behaviors (Nguyen et al., 2013; Rozière et al., 2020; Pan et al., 2024).
 With the rapid evolution of PLs and the diverse requirements of software applications, code translation has gained significant attention in both research and industry due to its wide range of practical applications.
@@
 It also facilitates efficient prototyping and development across multiple PLs for multi-platform software (Liu et al., 2023; Nguyen et al., 2015; Wu et al., 2010).
-However, achieving accurate automated code translation remains a challenging task due to the inherent differences in syntax and features among various PLs (Pan et al., 2024).
-Over the past few decades, various approaches have been proposed to automate code translation and address the associated challenges (Chen et al., 2025; Jiao et al., 2023).
+However, achieving accurate automated code translation remains a challenging task due to the inherent differences in syntax, features, and APIs across various PLs (Pan et al., 2024).
+Over the past few decades, various approaches have been proposed to automate code translation and address the associated challenges (Chen et al., 2025a; Jiao et al., 2023).
 Earlier methods primarily relied on statistical or neural machine translation (Nguyen et al., 2013; Chen et al., 2018; Rozière et al., 2020, 2022; Liu et al., 2023), while recent practices start leveraging the powerful large language models (LLMs) as code translators (Pan et al., 2024; Yang et al., 2024).
-Nevertheless, existing works mainly use a direct code-to-code translation approach, where the original program is taken as input to generate the translated program in a single step.
-In this step, LLMs may perform multiple tasks implicitly, such as understanding the semantics of the original program and generating the target program in another PL.
-Studies have shown that such direct code-to-code translation remains challenging, especially across PLs with significant differences in syntax and features (Pan et al., 2024; Macedo et al., 2025).
+However, the state-of-the-art practices with LLMs mainly use a direct code-to-code translation approach, where the original program is taken as input to generate the translated program in a single step (Yang et al., 2024).
+In this manner, LLMs may perform multiple tasks implicitly, such as understanding the semantics of the original program and implementing the semantics into the target program in another PL.
+Studies have shown that such direct translation remains challenging, notably across PLs with significant differences (Pan et al., 2024; Macedo et al., 2025).
 Decomposing challenging tasks into intermediate steps has proven effective in guiding LLMs to emulate successful human workflows in various coding tasks (e.g., splitting planning and implementation for code generation (Dong et al., 2024; Li et al., 2025) and fault localization and patch generation for debugging (Xia et al., 2024)).
+However, there are few experiences about whether decomposition also helps generate accurate code translations.
 Notably, we observe that code translation can also benefit from explicitly implementing intermediate steps that emulate human practices in translation.
-Specifically, human translators often perform semantic translation (Newmark, 1981) based on the meaning of the given text, first distilling the intent into a language-agnostic meaning and then rendering it in the target language (Newmark, 1981).
-Meanwhile, a recent study demonstrates that pseudocode, as a PL-agnostic representation of code intent and logic widely used in textbooks and research papers, can effectively guide code generation across PLs (Wu et al., 2025).
-Based on these insights, we further recognize that emulating semantic translation for code based on pseudocode can be a solution for the tasks that direct translation struggles with.
-For example, when translating a simple C++ program (Figure 1(a)) to Rust, all five studied LLMs (e.g., Qwen2.5-Coder-32B-Instruct) failed to produce an accurate translation via a one-step direct translation (Figure 1(b)).
-In comparison, following a clear pseudocode (Figure 1(c)) summarized from the original C++ program, these LLMs can successfully implement the original program’s functionality in Rust (Figure 1(d)).
-However, introducing explicit intermediate steps also has limitations.
-For example, it may introduce noise or suffer from semantic loss during information transmission (Liu et al., [n.
-d.]), which may conversely hinder accurate code translation.
-In general, the effectiveness, advantages, and limitations of such pseudocode-based code translation compared to the commonly adopted direct translation remain unclear.
-In this work, we conduct the first empirical study on pseudocode-based code translation to bridge the gap in understanding this under-explored approach.
+Specifically, human translators often perform semantic translation (Newmark, 1981) based on the semantics of the given text, first interpreting its meaning and then rendering the meaning in the target language (Newmark, 1981).
+In addition, a recent study demonstrates that pseudocode, as a PL-agnostic representation of code intent and logic widely used in textbooks and research papers, can effectively guide code generation across PLs (Wu et al., 2025).
+Based on these insights, we realize that emulating semantic translation for code based on pseudocode can be a solution for the tasks that direct translation struggles with.
+For example, when translating a simple C++ program (Figure 1(a)) to Rust, all five studied LLMs failed to produce an accurate translation via a one-step direct translation (Figure 1(b)).
+Meanwhile, following a clear pseudocode (Figure 1(c)) summarized from the original C++ program, these LLMs can successfully implement the original program’s functionality in Rust (Figure 1(d)).
+Nevertheless, the effectiveness, advantages, and limitations of the promising pseudocode-based code translation approach compared to the commonly adopted direct translation remain unclear.
+In this work, we conduct the first systematic empirical study on pseudocode-based code translation to bridge the gap in understanding this under-explored approach.
 Specifically, we assessed the performance of five translation strategies: the widely-adopted direct translation approach, two pseudocode-based strategies, and two hybrid strategies that combine direct and pseudocode-based translation.
-Through four research questions (RQs), we investigate the overall effectiveness of pseudocode-based code translation compared with the widely-adopted direct translation (RQ1), the helpfulness of pseudocode for varying PL pairs (RQ2), the advantages of pseudocode as an intermediary compared to a concrete PL, which also proved helpful in code translation (Macedo et al., 2025) (RQ3), and how the quality of pseudocode affects translation outcomes (RQ4).
+Through four research questions (RQs), we investigate the overall effectiveness of pseudocode-based code translation compared with the widely-adopted direct translation (RQ1), its helpfulness for varying PL pairs (RQ2), the advantages of pseudocode as an intermediary compared to a concrete PL, which also proved helpful in code translation (Macedo et al., 2025) (RQ3), and the impact of pseudocode quality on translation performance (RQ4).
 We also examine the successes and failures of pseudocode-based code translation via case studies and discuss future directions for further harnessing its potential benefits.
-The experiment is conducted with six popular PLs on solution programs of 323 LeetCode problems in three difficulty levels in LiveCodeBench (Jain et al., 2025), resulting in 9,690 translation tasks for each of five studied LLMs.
+The study is conducted with solution programs in six popular PLs for 323 LeetCode problems in three difficulty levels in LiveCodeBench (Jain et al., 2025), resulting in 9,690 translation tasks (i.e., 6 source PLs $\times$ 5 target PLs $\times$ 323 problems), for five popular LLMs.
 Our experimental results reveal several interesting findings and actionable insights.
-Specifically, we identify the complementary role of pseudocode-based code translation in effectively handling code translation tasks that direct translation struggles with.
-By combining the results of direct translation and pseudocode-based translation, the studied LLMs achieve average improvements in pass@10 rates over direct translation alone of 4.10%, 7.44%, and 13.75% on easy-, medium-, and hard-level problems, respectively.
-We also find that pseudocode-based translation can benefit varying source-target PL pairs, with more substantial helpfulness for translation from flexible PLs (e.g., Python) to more rigid PLs (e.g., Go) or involving the low-resource PL Rust.
-These findings suggest the adoption of a hybrid strategy combining direct and pseudocode-based translation results with further test-based selection to leverage the complementary advantages of both approaches, in particular when translating from flexible to rigid PLs or dealing with low-resource PLs.
-Moreover, with higher-quality pseudocode, we observed that the studied LLMs show the potential to achieve pass rates of 0.9646–0.9835, 0.8861–0.9512, and 0.6747–0.8286 on three-level tasks, respectively.
-The identified bottlenecks of pseudocode-based code translation stem from both code understanding and generation capabilities of LLMs.
-Our case studies also highlight the advantages of pseudocode-based translation in disentangling understanding and generation on complicated programs and mitigating distractions from the PL-specific details in original programs, as well as its limitations caused by incorrect, incomplete, or ambiguous pseudocode.
-These findings further illuminate the effective use of pseudocode-based translation and inspire future research on enhancing code translation through pseudocode.
-Figure 1. An Example of Correct C++-to-Rust Translation by Qwen2.5-Coder-32B-Instruct based on Pseudocode (The other four studied LLMs show similar symptoms in this case.)
+Specifically, we identify the complementary role of pseudocode-based code translation in effectively handling translation tasks that direct translation struggles with.
+By combining the results of direct translation and pseudocode-based translation, the studied LLMs achieve average improvements in translation accuracy over direct translation alone by 4.10%, 7.44%, and 13.76% on easy-, medium-, and hard-level problems, respectively.
+We also demonstrate that pseudocode-based translation can benefit various source-target PL pairs, with more significant benefits for translation from flexible PLs (e.g., Python) to more rigid PLs (e.g., Go), or involving Rust PL that has limited training resources.
+These findings suggest the adoption of a hybrid strategy that combines direct and pseudocode-based translation results with further test-based selection to leverage the complementary strengths of both approaches, particularly when handling the highly benefited PLs/PL pairs.
+Moreover, with higher-quality pseudocode, we observed that the studied LLMs can achieve pass rates of 0.9646–0.9835, 0.8861–0.9512, and 0.6747–0.8286 on three-level tasks, respectively, identifying bottlenecks of pseudocode-based code translation in both code understanding and code generation capabilities of LLMs.
+Our case studies further illustrate the advantages of pseudocode-based translation in decoupling the code understanding and generation burdens on complicated programs and mitigating distractions from the PL-specific details in the original programs, as well as its limitations caused by incorrect, incomplete, or ambiguous pseudocode.
+These findings illuminate the effective adoption of pseudocode-based translation to enhance LLMs’ code translation accuracy and provide evidence for future research on improving LLM-driven code translation.
+Figure 1.
+An Example of Correct C++-to-Rust Translation by Qwen2.5-Coder-32B-Instruct based on Pseudocode, where Direct Translation Failed.
+(The other four studied LLMs show similar symptoms in this case.)
 To summarize, this work makes the following contributions:
-  • We conduct the first empirical study on pseudocode-based code translation, exploring the effectiveness, advantages, and limitations of explicit semantic code translation via pseudocode in improving code translation accuracy by emulating human translation practices.
-  • We compare four pseudocode-based translation strategies with direct translation across 9,690 translation tasks based on programs in six popular PLs.
-With five popular LLMs as the code translator, we systematically investigate the effectiveness of pseudocode-based code translation in enhancing direct translation, its helpfulness for varying PL pairs, and its advantages and limitations, aiming to provide insights for its effective use and inspire future enhancement.
-  • Our study reveals interesting findings, including the complementary role of pseudocode-based translation in enhancing direct code translation, particularly for flexible-to-rigid PL translation and for low-resource Rust, and the potential and bottlenecks of pseudocode-based code translation.
-These findings yield actionable insights, including the adoption of a hybrid strategy combining direct and pseudocode-based translation results to enhance code translation accuracy, as well as the space in pseudocode quality to further harness the benefits of the approach.
+  • We conduct the first systematic empirical study on pseudocode-based code translation, exploring the effectiveness, advantages, and limitations of emulating human semantic code translation practices via pseudocode in improving code translation accuracy of LLMs.
+  • We compare four pseudocode-based translation strategies with direct translation across 9,690 translation tasks involving six popular PLs.
+With five popular LLMs as the translator, we systematically investigate the effectiveness of pseudocode-based code translation in enhancing direct translation, its helpfulness for varying PL pairs, and its advantages and limitations, aiming to provide insights for its effective adoption and evidence for its future enhancement.
+  • Our study reveals intriguing findings, including the complementary role of pseudocode-based translation in enhancing direct code translation, the highly benefited PLs/PL pairs, and the potential and bottlenecks of pseudocode-based code translation.
+These findings yield actionable insights, including the adoption of a hybrid strategy that combines direct and pseudocode-based translation results to enhance code translation accuracy, as well as identify improvement space in the generation of high-quality pseudocode to further harness the benefits of the approach.
+We release our experimental data, including 1,938 solution programs in six popular PLs collected for 323 LeetCode problems in LiveCodeBench-v5 to translate, as well as scripts to conduct direct and pseudocode-based translation and evaluate multilingual translation results.
+We will open-source them in our artifact (Artifact of this paper, ) to facilitate future research.
 The rest of the paper is organized as follows.
 Section 2 introduces the background of code translation and motivating examples.
-Section 3 presents the design of our empirical study, including the research questions, studied translation strategies, translation tasks, LLMs, and evaluation metrics.
+Section 3 presents the design of our empirical study, including the research questions, studied code translation strategies, tasks, LLMs, and evaluation metrics.
 Section 4 analyzes the experimental results to answer our research questions in detail and discusses the advantages and limitations of pseudocode-based code translation through case studies.
-Section 6 discusses related work.
+Section 5 and Section 6 discuss the threats to validity and related work, respectively.
 Finally, Section 7 concludes the paper.
@@
 2.1. Code Translation
-Code translation refers to the process of converting programs written in one programming language (PL) into another PL while preserving code semantics (Nguyen et al., 2013; Rozière et al., 2020; Pan et al., 2024).
+Code translation refers to the process of converting programs written in a programming language (PL) into another while preserving their code semantics (Nguyen et al., 2013; Rozière et al., 2020; Pan et al., 2024).
 Formally, given an original program $S=\langle s_{1},s_{2},\dots,s_{n}\rangle$ written in the source PL $L_{s}$, code translation aims to produce a translated program $T=\langle t_{1},t_{2},\dots,t_{m}\rangle$ in the target PL $L_{t}$, such that $T$ performs the same functionality as $S$, where $s_{i}$ and $t_{j}$ represent statements in $L_{s}$ and $L_{t}$, respectively.
 For example, we can translate the C++ program in Figure 1(a) and the Python program in Figure 2(a) to their semantically equivalent Rust program in Figure 1(d) and Java program in Figure 2(d), respectively.
-In code translation, there may not always be equivalent constructs, statements, and APIs in the target PL for those in the source PL, and vice versa, which makes code translation challenging.
+In code translation, there may not always be equivalent constructs, statements, and APIs in the target PL for those in the source PL, and vice versa, which makes code translation challenging (Yang et al., 2024; Macedo et al., 2025).
 In the past decade, automated code translation approaches have developed from traditional rule-based and statistical machine translation (Nguyen et al., 2013) to neural-model-based methods to automatically learn diverse and complex translation patterns across PLs (Chen et al., 2018; Rozière et al., 2020, 2022; Liu et al., 2023; Jiao et al., 2023).
 Recently, LLMs pre-trained on massive code corpora demonstrate superior capabilities in code translation (Yang et al., 2024; Ibrahimzada et al., 2025; Pan et al., 2024).
-These works mainly focus on a code-to-code translation approach, i.e., directly converting the original program in $L_{s}$ to the translated program in $L_{t}$ in an integrated step.
+These works mainly focus on a direct one-step code-to-code translation approach, i.e., directly converting the original program in $L_{s}$ to the translated program in $L_{t}$ in an integrated step.
 However, LLMs may struggle to precisely capture and replicate the semantics of the original program in another PL in a single step, leading to incorrect translations (Pan et al., 2024).
 To mitigate the gap between source and target PLs, Macedo et al.
-(2025) explored transitive translation via an intermediate PL, but they still focus on code-to-code translation and fail to handle certain translation tasks.
-In this study, we explore an alternative code translation approach by emulating human semantic translation in natural language translation (Newmark, 1981), which explicitly conducts code understanding by generating pseudocode as an intermediate step to interpret code intent and logic to facilitate translation.
-In code translation, the existence of original programs provides a reference to validate the translated programs.
-Unlike code generation tasks that rely on manually defined test cases or human evaluation to assess the generated code (Chen et al., 2021; Austin et al., 2021), code translation allows automated correctness assessment by comparing the outputs of the translated and original programs on the same test cases.
-These enable multiple attempts to be a feasible and meaningful practice to enhance code translation accuracy (Macedo et al., 2025; Zheng et al., 2023).
-In this study, we explore the translation setup of 10 attempts, which is a typical cost-effective setup that balances cost and gain (Macedo et al., 2025; Wu et al., 2025).
-It also enables the investigation of combining results from both direct and pseudocode-based code translation (Section 3.3).
+(2025) explored transitive translation via an intermediate PL, i.e., chaining two code-to-code translations.
+But they still fail to handle certain translation tasks.
+In this study, we explore an alternative code translation approach by emulating human semantic translation in natural language translation (Newmark, 1981), which explicitly conducts code interpretation by generating pseudocode as an intermediate step to understand code intent and logic to facilitate translation.
+A property of the code translation task is that the original program is available, enabling automatic correctness assessment of multiple translation candidates.
+Specifically, unlike the code generation tasks where the result correctness is typically evaluated using manually prepared test assertions or human judgment (Chen et al., 2021; Austin et al., 2021), code translation allows automatic correctness assessment of the translation results by comparing the outputs of the translated and original programs on the same test inputs.
+This property makes it practical and meaningful to generate multiple translation candidates to increase the likelihood of getting a correctly translated program, since the original program can help identify the best candidate as the final result via automatic correctness assessment (Rozière et al., 2020; Macedo et al., 2025; Zheng et al., 2023).
+In this study, we adopt ten generation attempts, which represents a common cost-effective trade-off between computational cost and performance gains on coding tasks (Macedo et al., 2025; Wu et al., 2025).
+The multiple-attempt setup also allows us to easily explore the effectiveness of combining results from both direct and pseudocode-based translation approaches (Section 3.3).
 2.2. Motivating Examples
 Figure 2. An Example of Correct Python-to-Java Translation by DeepSeek-Coder-V2-Lite based on Pseudocode
-To guide LLMs to solve complicated coding tasks, researchers often decompose a complex task into multiple simpler sub-tasks following human experience (e.g., problem solving/planning and implementation for code generation (Dong et al., 2024; Li et al., 2025; Wu et al., 2025)).
-Inspired by this, we explore whether decomposing code translation into sub-tasks following human practices can also facilitate LLMs to produce more accurate translations by emulating the successful human workflow.
-Recalling the practices in natural language translation, humans typically employ two approaches.
-When expressions or sentence structures in the source and target languages are sufficiently similar, a literal word-by-word translation is often adequate.
-However, when literal translation is challenged by significant differences in grammar, idiomatic usage, or context between the source and target languages, humans tend to adopt semantic translation (Newmark, 1981), where they first interpret the underlying meaning of the source sentence and then render it in the target language in the manner best to convey the intended message.
+To guide LLMs to solve complicated coding tasks, researchers propose to decompose a complex task into multiple simpler sub-tasks following human experience (e.g., problem solving/planning and implementation for code generation (Dong et al., 2024; Li et al., 2025; Wu et al., 2025)).
+Inspired by this, we explore whether decomposing code translation into sub-tasks following human practices can also facilitate LLMs to produce more accurate code translations by emulating the successful human workflow.
+Recall the human practices in natural language translation.
+When the expressions or sentence structures in source and target languages are sufficiently similar, a literal word-by-word translation is often feasible.
+Meanwhile, when human translators struggle with the significant differences in grammar, idiomatic usage, or context between source and target languages, humans tend to adopt semantic translation (Newmark, 1981), where they first interpret the underlying meaning of the source sentence and then render it in the target language in the manner best to convey the intended message.
 Interestingly, we observe that semantic interpretation also benefits LLMs in code translation.
 Specifically, we prompt LLMs to explicitly conduct semantic translation via pseudocode-based transitive translation, i.e., first producing a semantic interpretation (pseudocode) for code intent and logic of the original program, and then generating the translated program based on pseudocode.
-We adopt pseudocode to represent semantics since it is a widely adopted form for describing code intent and logic of algorithms in textbooks and research papers.
-It also proves effective in guiding LLMs to replicate the code intent and logic in varying PLs and can be synthesized from programs (Wu et al., 2025).
+We use pseudocode to represent semantics because it is a common form for describing code intent and logic of algorithms in textbooks and research papers.
+Also, it proves effective in guiding LLMs to replicate the code intent and logic in varying PLs and can be synthesized from programs (Wu et al., 2025).
 Figure 2 shows an example where pseudocode enables DeepSeek-Coder-V2-Lite (abbr.
@@
 Notably, all five studied LLMs (listed in Table 2) suffer from similar issues, and all of them except DSCoder generate a correct implementation when provided with pseudocode summarized by themselves as an intermediate step.
-These motivating examples suggest that incorporating pseudocode as a semantic representation for transitive translation can promote accurate code translations, mirroring the benefits of humans using semantic translation for natural languages.
+These motivating examples suggest that incorporating pseudocode as a semantic representation and conducting transitive translation can promote accurate code translations, mirroring the benefits of humans using semantic translation for natural languages.
 This is reasonable as the approach guides LLMs to understand code intent and logic and then perform code generation, decomposing a complex task into simpler sub-tasks, and LLMs are skillful in both tasks (Geng et al., 2024; Jain et al., 2025).
 This enables a new access to achieving code translation, which may bypass LLMs’ struggle with direct translation.
-Despite these benefits, the proper use of pseudocode in code translation remains underexplored, e.g., how to effectively translate code via pseudocode, how effective it is when handling distinct source-target PL pairs, and what strengths it entails.
-In addition, pseudocode-based code translation may also face challenges, e.g., the potential error propagation from incorrect pseudocode and the inherent ambiguity in natural language.
-The limitations it entails and potential remedies to overcome the limitations also remain unclear.
+Despite these benefits, the proper use of pseudocode in code translation remains underexplored, e.g., how to effectively translate code via pseudocode, how effective it is when handling different source-target PL pairs, and what strengths it entails.
+In addition, pseudocode-based code translation may also face challenges, e.g., the potential error propagation from incorrect pseudocode and the inherent ambiguity in natural language (Liu et al., ).
+The limitations it entails and potential remedies to overcome the limitations also remain unknown.
 To bridge these gaps, we conduct a systematic investigation of pseudocode-guided code translation.
-Our study aims to provide deeper insights into when and how pseudocode can effectively enhance code translation.
+Our study aims to provide insights into how emulating semantic translation via pseudocode can effectively enhance code translation.
 3. Study Design
@@
 RQ1: How effective is pseudocode-based code translation?
-This RQ aims to show a general picture of the effectiveness of semantic translation for code via pseudocode.
-We also investigate how such pseudocode-based code translation can work with the widely adopted direct code translation, e.g., fully replacing direct translation or complementing it by adopting both strategies simultaneously.
-RQ2: How effective is pseudocode-based semantic translation for different pairs of source and target programming languages?
+This RQ aims to show a general picture of the effectiveness of emulating semantic translation for code via pseudocode.
+We investigate whether pseudocode-based translation effectively complements direct translation and whether it can fully replace direct translation, aiming to shed light on the effective strategy to leverage translation approaches.
+RQ2: How helpful is pseudocode-based code translation for different pairs of source and target programming languages?
 Existing studies show that the code translation accuracy of LLMs varies across different pairs of source and target PLs, indicating that distinct PL pairs may pose varying challenges for LLMs (Pan et al., 2024; Macedo et al., 2025).
@@
 It also investigates whether certain PL pairs can benefit more from pseudocode-based translation, suggesting scenarios where pseudocode-based translation is highly recommended.
-RQ3: Is pseudocode a more effective intermediary compared to a specific programming language?
+RQ3: How effective is pseudocode as an intermediary compared to a specific programming language?
 Macedo et al.
-(2025) demonstrates that a specific PL with close syntactic and semantic similarities to the source and target PLs may act as an effective transitive intermediary to benefit LLMs in code translation.
-For example, translating Python code to Java via Rust is found to lead to better accuracy than direct translation.
-In this RQ, we aim to understand whether pseudocode can serve as a more effective intermediary than a specific transitive PL.
+(2025) demonstrates that a specific PL with close syntactic and semantic similarities to the source and target PLs also benefits transitive code translation.
+For example, translating Python code to Java via a Rust implementation as a bridge is found to complement direct Python-to-Java translation.
+In this RQ, we aim to understand whether pseudocode, which interprets mainly code intent and logic without PL-specific details, can serve as a more effective intermediary than a specific PL in transitive code translation.
 RQ4: How does the quality of pseudocode affect the effectiveness of pseudocode-based code translation?
-LLMs may not accurately interpret and record the intent of the input program, resulting in low-quality pseudocode that may conversely hinder LLMs from generating correct code translation.
-This RQ investigates the code translation performance based on high-quality pseudocode, aiming to reveal the potential of pseudocode-based code translation and identify the bottlenecks of studied LLMs to harness the potential effectiveness.
+LLMs may not accurately interpret the intent and logic of original programs, which may result in low-quality pseudocode, conversely hindering LLMs from generating correctly translated programs.
+This RQ investigates the code translation performance based on high-quality pseudocode, aiming to reveal the potential of pseudocode-based code translation and identify the bottlenecks of studied LLMs to harness the potential effectiveness of pseudocode-based code translation.
 3.2. Experimental Translation Subjects
@@
 In this study, we evaluate code translation performance across six widely studied PLs, i.e., C++, Python, Java, JavaScript, Go, and Rust.
-The choice of these PLs follows prior work, where translation among the first four PLs has been extensively explored (Rozière et al., 2022; Jiao et al., 2023; Yang et al., 2024), while recent studies have highlighted the emerging PLs such as Go and Rust (Pan et al., 2024; Macedo et al., 2025).
+The choice of these PLs follows prior work, where translation among the first four popular PLs (TIOBE Index, ) has been extensively explored (Rozière et al., 2022; Jiao et al., 2023; Yang et al., 2024), while recent studies have highlighted the emerging Go and Rust (Pan et al., 2024; Macedo et al., 2025).
 3.2.2. Code Translation Tasks
-We prompt LLMs to translate the solution programs of 323 LeetCode problems in LiveCodeBench (Jain et al., 2025).
-We choose to translate such programs since they involve common implementations widely used in practical development and become a representative benchmark to evaluate the coding ability of LLMs (Hui et al., 2024; Team, 2024).
-Actually, solution programs of LeetCode problems have been widely adopted as the subjects in code translation studies since they formulate easy-to-collect multilingual parallel programs (Wang et al., 2023a; Ahmad et al., 2023).
-Moreover, LiveCodeBench provides carefully validated and representative tests, facilitating the reliable validation of the translation results.
+We prompt LLMs to translate the solution programs of 323 LeetCode problems in LiveCodeBench-v5 (Jain et al., 2025).
+We choose these programs because they involve common implementations widely used in practical development and form a representative benchmark to evaluate the coding ability of LLMs (Hui et al., 2024; Team, 2024).
+Actually, solution programs of LeetCode problems have been widely adopted as subjects in code translation studies, as they provide easy-to-collect multilingual parallel programs (Wang et al., 2023a; Ahmad et al., 2023).
+Besides, LiveCodeBench offers carefully validated and representative tests, which facilitate the validation of translation results.
 Additionally, the LiveCodeBench programming problems are released after May 2023 (Jain et al., 2025), allowing the evaluation to be less subject to the data contamination issue compared to the conventional code translation benchmarks like CodeNet (Puri et al., 2021), TransCoder (Rozière et al., 2020), and AVATAR (Ahmad et al., 2023) established in 2020–2021.
-To support the investigation of code translation among diverse PLs, we follow existing practices (Jain et al., 2025; Wu et al., 2025) to collect correct user-written solution programs that can pass all the tests on LiveCodeBench, for all six PLs.
-These solution programs are used as the source programs to translate.
-Among all 381 LeetCode problems from LiveCodeBench, we excluded 16 problems with incorrect tests identified by Wu et al.
+Table 1. AST Node Counts and API Invocation Counts of Programs for Different Difficulty-Level Problems
+Difficulty Average Count of AST Nodes Average Count of Invocation Nodes
+C++ Go Java JS Python Rust C++ Go Java JS Python Rust
+Easy
+167.91
+142.81
+159.33
+138.50
+111.22
+173.58
+2.56
+2.29
+2.94
+1.97
+3.32
+5.16
+Medium
+280.25
+263.34
+267.83
+266.81
+188.18
+298.54
+5.43
+4.95
+5.46
+4.42
+5.64
+8.69
+Hard
+592.81
+529.85
+519.15
+546.64
+339.70
+584.04
+10.55
+8.85
+7.43
+11.23
+8.77
+13.68
+To support the investigation of code translation among diverse PLs, we follow existing practices (Jain et al., 2025; Wu et al., 2025) to collect user-written solution programs that can pass all the tests on LiveCodeBench, for all six PLs.
+These solution programs are used as the original programs to translate.
+Among all 381 LeetCode problems from LiveCodeBench-v5, we excluded 16 problems with incorrect tests identified by Wu et al.
 (2025).
@@
 This results in 116 easy problems, 160 medium problems, and 47 hard problems.
+The solution programs of harder problems typically include more complex solving logic and implementations (Jain et al., 2025).
+As shown in Table 1, the average abstract syntax tree (AST) node count and API invocation count [^1] The statistics are based on counts of all AST nodes and external function calls identified by tree-sitter (Developers, ), respectively.
+of solution programs for medium/hard-level problems are 1.7x–1.9x / 3.1x–3.9x and 1.7x–2.2x / 2.5x–5.7x as those of easy-level problems across six PLs, suggesting that the harder tasks present higher translation complexity (Ou et al., 2024).
 Note that, for each problem, LiveCodeBench does not provide corresponding solution programs and only supports the evaluation of Python solutions.
-We collect the solution programs in six PLs and extend the evaluation scaffold to all six studied PLs via execution-based validation, following Wu et al.
-(2025) (who extended the support of C++ and Rust).
-We open-source these materials to facilitate future research on code translation (Artifact of this paper, [n.
-d.]).
+We collect the solution programs in six PLs and extend the execution-based evaluation scaffold to all six studied PLs.
 […]
 3.3. Experimental Strategies for Code Translation
+(a) Prompt for Direct Translation
+(b) Prompt for Generating Code from Pseudocode
+(c) Prompt for Generating Code from Pseudocode with Original Program as Reference Implementation
+Figure 3. Prompts Used in the Study (The examples are instantiated with the translation tasks in Figure 1).
 In this study, we investigate the effectiveness of five prompting strategies for code translation.
@@
     Direct Translation (abbr.
-$\mathcal{D}$): This strategy prompts LLMs to translate the programs in the source PL to the target PL.
-The strategy is widely adopted by existing LLM-based code translation studies (Yang et al., 2024; Pan et al., 2024).
+$\mathcal{D}$): This strategy prompts LLMs to translate original programs in the source PL to the target PL in one step.
+The strategy is widely used by existing LLM-based code translation studies (Yang et al., 2024; Pan et al., 2024).
 We follow these studies to design the prompt as shown in Figure 3(a).
@@
     Transitive Translation via Pseudocode (abbr.
-$\mathcal{P}$): This strategy prompts LLMs to explicitly conduct semantic translation using two transitive steps: (i) generate a piece of pseudocode based on the original program following the prompt from Wu et al.
-(2025); and (ii) implement the pseudocode into a translated program in the target PL following the prompt in Figure 3(b).
+$\mathcal{P}$): This strategy prompts LLMs to emulate semantic translation using two transitive steps explicitly: (i) generate a piece of pseudocode to describe the PL-agnostic code intent and logic for the original program; and (ii) implement the pseudocode into a program in the target PL following the prompt in Figure 3(b).
 As introduced in Section 2.2, this strategy is inspired by human practices of translation and the findings that pseudocode can encode the PL-agnostic code intent and logic of a program (Wu et al., 2025), serving as an appropriate semantic representation for emulating semantic translation.
 In our study, pseudocode and translated programs are generated by the same LLM (referred to as “translator LLM”).
-The prompts are designed following the practices of Wu et al.
+Our prompts are designed following practices of Wu et al.
 (2025).
+In particular, we reuse their prompt for pseudocode generation, which guides LLMs to generate pseudocode following: 1) using indents to represent control structures; 2) avoiding usages and type information specific to PLs; 3) omitting implementation of common algorithms, data structures, and operations; and 4) naming functions using the template “function [FUNC_NAME] ([PARAM_LIST])” (e.g., “function max(a, b)”).
+These conventions align with the general expectations of pseudocode style in prior studies (Kulal et al., 2019; Kulkarni et al., ; Acharjee et al., 2022).
+Given that there is no unified format and grammar for pseudocode, we allow LLMs to generate pseudocode in their most familiar style based on these conventions.
+This prompt is available in our artifact (Artifact of this paper, ).
   3. (3)
     Transitive Translation via Pseudocode with Original Program as Context (abbr.
-$\mathcal{PC}$): This strategy extends the above pseudocode-based strategy to include more implementation details (i.e., original program) as the context to guide code translation.
+$\mathcal{PC}$): This strategy extends the above pseudocode-based strategy to include implementation details (i.e., original program) as the context to guide code translation.
 Specifically, considering that pseudocode abstracts implementation details of the original program, it may miss some helpful information for LLMs to implement the original program’s intent and logic in the target PL.
-Thus, we include the original program in the source PL as an implementation example of the pseudocode for LLMs’ reference when generating the target code through explicit semantic translation, using the prompt shown in Figure 3(c).
+Thus, we include the original program in the source PL as an implementation example of the pseudocode for LLMs’ reference when generating the translated program by emulating semantic translation, using the prompt shown in Figure 3(c).
 As mentioned in Section 2.1, we generate ten translations for each original program. When using the above strategies, we repeatedly generate ten translations using their respective prompts.
-We also examine the effectiveness of combining direct translation with pseudocode-based translation.
-Specifically, we design two straightforward hybrid strategies by evenly mixing translation results of direct translation and transitive translation via pseudocode, while keeping the same total attempts as the above strategies (i.e., 10 attempts as mentioned in Section 2.1, with five attempts for each approach).
+We also examine the effectiveness of combining direct translation and pseudocode-based translation.
+Specifically, we design two straightforward hybrid strategies by evenly mixing direct translation results and pseudocode-based translation results.
+We keep the same total attempts of ten as the above three single-approach strategies, thus combining five translated programs generated by each approach.
 The two hybrid strategies are defined as follows:
@@
 We include these hybrid strategies to study whether combining direct translation and pseudocode-based translation can leverage their complementary strengths across different translation tasks.
-(a) Prompt for Direct Translation
-(b) Prompt for Generating Code from Pseudocode
-(c) Prompt for Generating Code from Pseudocode with Source Program as Reference Implementation
-Figure 3. Prompts Used in the Study (The examples are instantiated with the translation tasks in Figure 1).
 3.4. Experimental Setups
@@
 The metric measures the ratio of tasks on which the translator LLM successfully generates a correct translation in at least one out of ten attempts.
-Compared with text-based metrics like BLEU and Exact Match rate, the execution-based pass@10 metric can measure the semantic equivalence of the translation results based on the high-quality tests provided in LiveCodeBench.
+Compared with text-based metrics like BLEU and Exact Match rate, the execution-based pass@10 metric can measure the semantic equivalence of the translation results based on the tests provided by LiveCodeBench.
 The metric is widely adopted in code generation (Jain et al., 2025; Wu et al., 2025) and translation studies (Macedo et al., 2025).
-Note that we use pass@10 rather than using pass@1 for one attempt, because we aim to investigate the maximum potential of the studied strategies and LLMs in code translation, considering the higher translation accuracy achieved with ten attempts than one single attempt demonstrated by existing study (Macedo et al., 2025), as well as the practical feasibility of filtering out the incorrect translations based on the source program’s behavior as discussed in Section 2.1.
-The result of pass@1 shows similar trends.
-Studied LLMs: In this study, we investigate five popular LLMs from different famous LLM families deployable on our machine.
+Note that we use pass@10 to investigate the potential of the studied translation strategies and LLMs, considering the higher translation accuracy achieved with ten attempts than one single attempt demonstrated by existing study (Macedo et al., 2025) and the practical feasibility of filtering translation results by referring to the original program’s behavior as we discuss in Section 2.1.
+In addition, we discuss pass@{2,4,6,8} in RQ1.
+Studied LLMs: In this study, we investigate five popular LLMs deployable on our machine.
 They are Qwen2.5-Coder-32B-Instruct (abbr.
 Qwen32B) (Hui et al., 2024), Qwen2.5-Coder-7B-Instruct (abbr.
-Qwen7B) (Hui et al., 2024), Phi-4 (Abdin et al., 2024), GPT-4o-mini (OpenAI, [n.
-d.]), and DeepSeek-Coder-V2-Lite-Instruct (abbr.
+Qwen7B) (Hui et al., 2024), Phi-4 (Abdin et al., 2024), GPT-4o-mini (OpenAI, ), and DeepSeek-Coder-V2-Lite-Instruct (abbr.
 DSCoder) (DeepSeek-AI, 2024).
@@
 (abbr. Qwen32B)
-Qwen Open Source 32.8B Mar 2024 (Knowledge Cutoff Information of Qwen32B, [n. d.])
-Phi-4 Microsoft Open Source 14.7B May 2024 (Knowledge Cutoff Information of Phi-4, [n. d.])
-GPT-4o-mini OpenAI Commercial (Unknown) Oct 2023 (Knowledge Cutoff Information of GPT-4o-mini, [n. d.])
+Qwen Open Source 32.8B Mar 2024 (Knowledge Cutoff Information of Qwen32B, )
+Phi-4 Microsoft Open Source 14.7B May 2024 (Knowledge Cutoff Information of Phi-4, )
+GPT-4o-mini OpenAI Commercial (Unknown) Oct 2023 (Knowledge Cutoff Information of GPT-4o-mini, )
 Qwen2.5-Coder-7B-Instruct
 (abbr. Qwen7B)
-Qwen Open Source 7.62B Mar 2024 (Knowledge Cutoff Information of Qwen7B, [n. d.])
+Qwen Open Source 7.62B Mar 2024 (Knowledge Cutoff Information of Qwen7B, )
 DeepSeek-Coder-V2-Lite
 -Instruct (abbr. DSCoder)
-DeepSeek Open Source 15.7B Nov 2023 (Knowledge Cutoff Information of DSCoder, [n. d.])
-*: We run Qwen2.5-Coder-32B-Instruct-GPTQ-Int4 on with NVIDIA RTX4090 GPU.
+DeepSeek Open Source 15.7B Nov 2023 (Knowledge Cutoff Information of DSCoder, )
+*: We run Qwen2.5-Coder-32B-Instruct-GPTQ-Int4 with NVIDIA RTX4090 GPU.
 The quantized GPTQ-Int4 version is found to show comparable coding ability as the original bf16 version while taking less GPU memory (Wu et al., 2025).
-LLM Configuration: We run the four open-source LLMs on our server and access the commercial GPT-4o-mini via OpenAI API.
-To balance the diversity and reliability, we adopt the temperature of 0.2 following the practice in code generation studies (Jain et al., 2025; Wu et al., 2025; Xu et al., 2024).
-The maximum output tokens are set to be 3000 (i.e., 1.5x maximum input code length) to accommodate normal code translation outputs.
+LLM Configuration: We run the four open-source LLMs on our machine and access the commercial GPT-4o-mini via OpenAI API.
+To balance the diversity and reliability, we adopt the temperature of 0.2 following the practice in LLM-based coding studies (Jain et al., 2025; Wu et al., 2025; Xu et al., 2024).
+The maximum output tokens are set to be 3000 (representing 1.5x maximum input code length) to accommodate normal translation outputs.
 The other configurations (e.g., top_p and penalty) are kept as default values.
 To mitigate randomness, we repeat the experiments three times and report the average results.
-Experimental Environment: We run experiments on our server with Ubuntu 22.04 OS.
-The server is equipped with NVIDIA RTX4090 GPUs to deploy the open-source LLMs, as well as an AMD Ryzen Threadripper PRO 3995WX 64-Core CPU to support the parallel evaluation of massive generated translation results.
+Experimental Environment: We run experiments on a machine with Ubuntu 22.04 OS.
+The server is equipped with NVIDIA RTX4090 GPUs to deploy the open-source LLMs, as well as an AMD Ryzen Threadripper PRO 3995WX 64-Core CPU to support the parallel evaluation of plenty of generated translation results.
 4. Results and Analysis
-4.1. RQ1: Overall Effectiveness of Pseudocode-based Transitive Code Translation
+4.1. RQ1: Overall Effectiveness of Pseudocode-based Code Translation
 In RQ1, we investigate the overall effectiveness of pseudocode-based code translation.
-Specifically, we explore (1) how pseudocode-based code translation can help enhance the performance of the current widely-adopted direct translation strategy; and (2) whether pseudocode generated by the translator LLM can effectively represent the code intent or logic of the original program under translation and promote accurate code translation results.
+Specifically, we explore (1) whether pseudocode-based code translation can advance the accuracy of the current widely-adopted direct translation approach, by either complementarily enhancing or completely replacing direct translation; and (2) whether pseudocode generated by the translator LLM can effectively represent the code intent or logic of the original program under translation and promote accurate code translation results.
 We investigate these by comparing the performance of five translation strategies introduced in Section 3.3.
@@
 Table 3 shows the pass@10 rate of each LLM using different strategies.
-Table 2. Pass@10 Rate of Code Translation with Different Translation Strategies
+Table 3. Pass@10 Rate of Code Translation Using Different Translation Strategies
 LLM Strategy Easy Tasks Medium Tasks Hard Tasks
@@
 $\mathcal{P}$  (viaPseudocode) 0.9442 (-0.40%) 0.8736 (-0.41%) 0.6768 (-4.12%)
-$\mathcal{PC}$  (viaPseudocode w/ program as context) 0.9648 (+1.77%) 0.9001 (+2.61%) 0.7553 (+7.00%)
-$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudo.]) 0.9773 (+3.09%) 0.9403 (+7.19%) 0.7948 (+12.59%)
-$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudo. w/ prog.]) 0.9773 (+3.09%) 0.9327 (+6.33%) 0.7988 (+13.16%)
+$\mathcal{PC}$  (viaPseudocode w/ original program as context) 0.9648 (+1.77%) 0.9001 (+2.61%) 0.7553 (+7.00%)
+$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudocode]) 0.9773 (+3.09%) 0.9403 (+7.19%) 0.7948 (+12.59%)
+$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudocode w/ context]) 0.9773 (+3.09%) 0.9327 (+6.33%) 0.7988 (+13.16%)
 Phi-4 $\mathcal{D}$  (Direct Translation) 0.9276 0.8313 0.6281
 $\mathcal{P}$  (viaPseudocode) 0.8782 (-5.33%) 0.7744 (-6.84%) 0.5832 (-7.15%)
-$\mathcal{PC}$  (viaPseudocode w/ program as context) 0.9254 (-0.24%) 0.8347 (+0.41%) 0.6463 (+2.90%)
-$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudo.]) 0.9646 (+3.99%) 0.8947 (+7.63%) 0.7083 (+12.77%)
-$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudo. w/ prog.]) 0.9610 (+3.60%) 0.8903 (+7.10%) 0.7038 (+12.05%)
+$\mathcal{PC}$  (viaPseudocode w/ original program as context) 0.9254 (-0.24%) 0.8347 (+0.41%) 0.6463 (+2.90%)
+$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudocode]) 0.9646 (+3.99%) 0.8947 (+7.63%) 0.7083 (+12.77%)
+$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudocode w/ context]) 0.9610 (+3.60%) 0.8903 (+7.10%) 0.7040 (+12.08%)
 GPT-4o -mini $\mathcal{D}$  (Direct Translation) 0.9279 0.8552 0.6667
 $\mathcal{P}$  (viaPseudocode) 0.8647 (-6.81%) 0.7219 (-15.59%) 0.5000 (-25.00%)
-$\mathcal{PC}$  (viaPseudocode w/ program as context) 0.9239 (-0.43%) 0.8635 (+0.97%) 0.6837 (+2.55%)
-$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudo.]) 0.9658 (+4.08%) 0.9054 (+5.87%) 0.7404 (+11.05%)
-$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudo. w/ prog.]) 0.9621 (+3.69%) 0.9144 (+6.92%) 0.7617 (+14.25%)
+$\mathcal{PC}$  (viaPseudocode w/ original program as context) 0.9239 (-0.43%) 0.8635 (+0.97%) 0.6837 (+2.55%)
+$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudocode]) 0.9658 (+4.08%) 0.9054 (+5.87%) 0.7404 (+11.05%)
+$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudocode w/ context]) 0.9621 (+3.69%) 0.9144 (+6.92%) 0.7617 (+14.25%)
 Qwen7B $\mathcal{D}$  (Direct Translation) 0.9073 0.7928 0.5400
 $\mathcal{P}$  (viaPseudocode) 0.7735 (-14.75%) 0.5928 (-25.23%) 0.3286 (-39.15%)
-$\mathcal{PC}$  (viaPseudocode w/ program as context) 0.8989 (-0.93%) 0.7929 (+0.01%) 0.5480 (+1.48%)
-$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudo.]) 0.9336 (+2.90%) 0.8395 (+5.89%) 0.5849 (+8.31%)
-$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudo. w/ prog.]) 0.9420 (+3.82%) 0.8568 (+8.07%) 0.6154 (+13.96%)
+$\mathcal{PC}$  (viaPseudocode w/ original program as context) 0.8989 (-0.93%) 0.7929 (+0.01%) 0.5480 (+1.48%)
+$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudocode]) 0.9336 (+2.90%) 0.8395 (+5.89%) 0.5849 (+8.31%)
+$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudocode w/ context]) 0.9420 (+3.82%) 0.8568 (+8.07%) 0.6154 (+13.96%)
 DSCoder $\mathcal{D}$  (Direct Translation) 0.8964 0.7876 0.5586
 $\mathcal{P}$  (viaPseudocode) 0.8562 (-4.48%) 0.7335 (-6.87%) 0.5069 (-9.26%)
-$\mathcal{PC}$  (viaPseudocode w/ program as context) 0.9218 (+2.83%) 0.7875 (-0.01%) 0.5530 (-1.00%)
-$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudo.]) 0.9482 (+5.78%) 0.8569 (+8.80%) 0.6442 (+15.32%)
-$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudo. w/ prog.]) 0.9537 (+6.39%) 0.8581 (+8.95%) 0.6459 (+15.63%)
+$\mathcal{PC}$  (viaPseudocode w/ original program as context) 0.9218 (+2.83%) 0.7875 (-0.01%) 0.5530 (-1.00%)
+$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudocode]) 0.9482 (+5.78%) 0.8569 (+8.80%) 0.6442 (+15.32%)
+$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudocode w/ context]) 0.9537 (+6.39%) 0.8581 (+8.95%) 0.6459 (+15.63%)
 (Average) $\mathcal{D}$  (Direct Translation) 0.9214 0.8288 0.6199
 $\mathcal{P}$  (viaPseudocode) 0.8634 (-6.30%) 0.7392 (-10.81%) 0.5191 (-16.26%)
-$\mathcal{PC}$  (viaPseudocode w/ program as context) 0.9270 (+0.60%) 0.8357 (+0.83%) 0.6373 (+2.81%)
-$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudo.]) 0.9579 (+3.96%) 0.8874 (+7.06%) 0.6945 (+12.04%)
-$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudo. w/ prog.]) 0.9592 (+4.10%) 0.8905 (+7.44%) 0.7051 (+13.75%)
+$\mathcal{PC}$  (viaPseudocode w/ original program as context) 0.9270 (+0.60%) 0.8357 (+0.83%) 0.6373 (+2.81%)
+$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudocode]) 0.9579 (+3.96%) 0.8874 (+7.06%) 0.6945 (+12.04%)
+$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudocode w/ context]) 0.9592 (+4.10%) 0.8905 (+7.44%) 0.7052 (+13.76%)
 Effectiveness in Enhancing Direct Translation.
-The results show that the two hybrid strategies (i.e., $\mathcal{D}\&\mathcal{P}$ and $\mathcal{D}\&\mathcal{PC}$) that combine results from pseudocode-based and direct translation effectively improved the code translation accuracy achieved by the baseline ($\mathcal{D}$).
+The results show that the two hybrid strategies (i.e., $\mathcal{D}\&\mathcal{P}$ and $\mathcal{D}\&\mathcal{PC}$) that combine translation outcomes from pseudocode-based and direct translation effectively improved the code translation accuracy achieved by the baseline ($\mathcal{D}$).
 Specifically, $\mathcal{D}\&\mathcal{P}$ improved the pass@10 rate by 3.96%, 7.06%, and 12.04% on average on three difficulty levels, respectively.
-$\mathcal{D}\&\mathcal{PC}$ led to more significant improvements of 4.10%, 7.44%, and 13.75% on three levels on average, respectively.
+$\mathcal{D}\&\mathcal{PC}$ led to more significant improvements of 4.10%, 7.44%, and 13.76% on three levels on average, respectively.
 The results demonstrate that pseudocode-based code translation enables LLMs to correctly translate many programs that cannot be correctly handled by direct translation alone.
@@
 The results indicate that pseudocode-based code translation can effectively complement the widely adopted direct translation to enhance code translation accuracy.
-Meanwhile, we observed that the two pure pseudocode-based translation strategies (i.e., $\mathcal{P}$ and $\mathcal{PC}$) only produce comparable or even worse performance than direct translation ($\mathcal{D}$).
+Meanwhile, we observed that the two pure pseudocode-based translation strategies (i.e., $\mathcal{P}$ and $\mathcal{PC}$) produced only comparable or even worse performance than direct translation ($\mathcal{D}$).
 Specifically, $\mathcal{PC}$ increased pass@10 of $\mathcal{D}$ by only 0.60%, 0.83%, and 2.81% on three levels on average; $\mathcal{P}$ even led to drops in pass@10 rate by 6.30%, 10.81%, and 16.26%.
-The results demonstrate that when conducting semantic translation via pseudocode, LLMs also failed to translate a few programs that could be correctly translated by direct translation, indicating that direct translation and pseudocode-based translation are complementary at the current stage.
-We discuss potential causes of this underperformance via case studies in Section 4.5; we also explore the potential of pure pseudocode-based code translation based on higher-quality pseudocode in RQ4 (Section 4.4).
+The results demonstrate that when conducting semantic translation via pseudocode, LLMs also failed to translate a few programs that could be correctly translated by direct translation.
+This indicates that direct translation and pseudocode-based translation are complementary at the current stage.
+We discuss how this underperformance can be attributed to incorrect, incomplete, and ambiguous pseudocode via case studies in Section 4.5, and we also explore the potential of pure pseudocode-based code translation based on higher-quality pseudocode in RQ4 (Section 4.4).
 Difficulty-wise.
-We also notice that the improvement from hybrid strategies is more significant for harder tasks.
-Specifically, the improvement in pass@10 rate brought by $\mathcal{D}\&\mathcal{P}$ and $\mathcal{D}\&\mathcal{PC}$ over $\mathcal{D}$ increases by 3.96% and 4.10% on easy-level tasks, while 12.04% and 13.75% on hard-level tasks on average, respectively.
-The results indicate that pseudocode-based code translation is notably helpful to complement direct translation on relatively complicated programs.
-Message 1: Pseudocode-based code translation can effectively complement the widely adopted direct translation strategy to translate programs that cannot be correctly handled by direct translation alone, with a more significant improvement for relatively complicated programs.
-We recommend practitioners adopt hybrid strategies that combine the strengths of direct translation and pseudocode-based translation to enhance code translation accuracy, by collecting translated programs from both approaches and identifying the final result based on tests.
+We also notice that the improvement of hybrid strategies is more significant on harder tasks.
+Specifically, the improvement in pass@10 rate brought by $\mathcal{D}\&\mathcal{P}$ and $\mathcal{D}\&\mathcal{PC}$ over $\mathcal{D}$ increases by 3.96% and 4.10% on easy-level tasks, while 12.04% and 13.76% on hard-level tasks on average, respectively.
+As introduced in Section 3.2, the original programs of harder LeetCode problems typically include more complex solving logic and implementations, involving more AST nodes and more external API invocations to translate.
+The results indicate that pseudocode-based code translation is notably helpful in complementing direct translation for relatively complicated programs.
+Our case studies in Section 4.5.1 will further discuss the advantages of pseudocode-based translation, which also align with this observation.
+Figure 4. Pass Rates of Direct Translation $\mathcal{D}$ and Hybrid Strategy $\mathcal{D}\&\mathcal{PC}$ with Different Numbers of Attempts
+[…]
+Message 1: Pseudocode-based code translation can effectively complement the widely adopted direct translation approach to translate programs that cannot be correctly handled by direct translation alone, with a more significant improvement for more complicated programs.
+We recommend that practitioners adopt hybrid strategies that combine the strengths of direct translation and pseudocode-based translation to enhance translation accuracy.
 Usefulness of Original Programs.
@@
 Specifically, the pure pseudocode-based strategy with the original program as context ($\mathcal{PC}$) achieved an average pass@10 rate of 0.9270, 0.8357, and 0.6373 on the three levels’ solution programs, respectively, while the strategy without the original program ($\mathcal{P}$) only achieved pass@10 rates of 0.8634, 0.7392, and 0.5191 accordingly.
-The results demonstrate that the pseudocode generated by the translator LLM missed or misinterpreted details necessary for the correct generation of programs in the target PL in some cases.
+The results suggest that the pseudocode generated by the translator LLM missed or misinterpreted details necessary for the correct generation of programs in the target PL in some cases.
 We investigate the concrete symptoms of this issue via case studies and discuss them in Section 4.5.
@@
 Specifically, the pass@10 rates of $\mathcal{D}\&\mathcal{P}$ and $\mathcal{D}\&\mathcal{PC}$ are generally comparable (with less than 0.01 difference in pass@10) in most comparisons.
-This indicates that the original program is less necessary to supplement pseudocode when using hybrid strategies, where direct translation may already provide access to referencing details in the original program for cases that highly rely on such details.
-The exceptional cases include GPT-4o-mini on hard-level tasks and Qwen7B on medium-level and hard-level tasks, where the hybrid strategy with the original program as context ($\mathcal{D}\&\mathcal{PC}$) achieved higher pass@10 rates.
-Therefore, it may still be beneficial to include the original program as a backup reference implementation of pseudocode when translating relatively complicated programs.
+This suggests that the original program is less necessary to supplement pseudocode when using hybrid strategies, as direct translation may already provide access to referencing details in the original program for translation tasks that heavily rely on such details, further confirming the complementary strengths of both approaches.
+The exceptional cases include GPT-4o-mini on hard-level tasks and Qwen7B on medium-level and hard-level tasks, where the hybrid strategy with the original program as context ($\mathcal{D}\&\mathcal{PC}$) achieved obviously higher pass@10 rates.
+Therefore, it is generally beneficial to include the original program as a backup reference implementation of pseudocode when translating relatively complicated programs.
 Message 2: The pseudocode generated by translator LLMs may miss details necessary for generating semantically equivalent programs in the target PL.
@@
 Differences in syntax and semantics among PLs may cause varying translation difficulty across PL pairs (Pan et al., 2024; Macedo et al., 2025).
-In this RQ, we thus further study the effectiveness of the pseudocode-based translation strategies on different PL pairs.
+In this RQ, we thus further compare the effectiveness of the pseudocode-based translation strategies on different source-target PL pairs.
 The comparison aims to (1) verify whether the helpfulness of pseudocode-based translation identified in RQ1 persists across different source and target PLs, and (2) explore if the pseudocode-based translation leads to significant improvements for certain PL pairs.
 The empirical findings are intended to guide developers in adopting pseudocode-based approaches for code translation involving different PLs.
-In this RQ, we focus on the improvement of the hybrid translation strategy with code ($\mathcal{D}\&\mathcal{PC}$), which exhibited generally optimal performance in RQ1, relative to the conventional direct translation strategy ($\mathcal{D}$).
-(a) Qwen-32B (Easy Tasks)
-(b) Qwen-32B (Med Tasks)
-(c) Qwen-32B (Hard Tasks)
+In this RQ, we focus on the improvement of the hybrid translation strategy with context ($\mathcal{D}\&\mathcal{PC}$), which showed generally optimal performance in RQ1, relative to the conventional direct translation strategy ($\mathcal{D}$).
+(a) Qwen32B (Easy Tasks)
+(b) Qwen32B (Med Tasks)
+(c) Qwen32B (Hard Tasks)
 (d) Phi-4 (Easy Tasks)
@@
 (i) GPT-4o-mini (Hard Tasks)
-(j) Qwen-7B (Easy Tasks)
-(k) Qwen-7B (Med Tasks)
-(l) Qwen-7B (Hard Tasks)
+(j) Qwen7B (Easy Tasks)
+(k) Qwen7B (Med Tasks)
+(l) Qwen7B (Hard Tasks)
 (m) DSCoder (Easy Tasks)
@@
 (o) DSCoder (Hard Tasks)
-Figure 4. Relative Improvement in Pass@10 Using “Direct + viaPseudocode (w/ code)” Strategy Compared to “Direct” Strategy for Code Translation. (The greener, the more improvement.)
+Figure 5.
+Relative Improvement in Pass@10 Using Hybrid Strategy $\mathcal{D}\&\mathcal{PC}$ Compared with Direct Translation Strategy $\mathcal{D}$ Across Different Source-Target Programming Languages.
+(The greener, the more improvement.)
 Overall Effectiveness.
@@
 Across all 450 PL-pair-model-difficulty combinations, 147 (32.7%) combinations achieve even more than 10% improvement.
-The improvement is consistently observed across different LLMs and task difficulties, with more significant improvement on harder tasks and weaker LLMs, as observed in RQ1.
+The improvement is consistently observed across different LLMs and task difficulties, with more significant improvement on harder tasks and weaker LLMs, similar to trends observed in RQ1.
 The results confirm the generalizable benefits of adopting pseudocode-based translation among different PL pairs.
@@
 As shown by the different colors in heatmaps within Figure 5, the improvement varies across different PL pairs, which may be due to the differences across PL pairs and the varying training corpus of different PLs.
-Nevertheless, we observe consistent trends across different LLMs and task difficulties.
-Specifically, we first noticed that the improvement of code translation from flexible and lightweight PLs (e.g., Python and JavaScript) to stricter and more complex PLs (e.g., Rust and Go) is usually more significant than the improvement in the opposite direction, as shown by the deeper green colors near the right-bottom corner of each heatmap.
+Nevertheless, we observe several consistent trends across different LLMs and task difficulties.
+Specifically, we first noticed that the improvement of code translation from flexible and lightweight PLs (e.g., Python and JavaScript) to stricter and more complex PLs (e.g., Rust and Go) is often more significant than the improvement in the opposite direction, as shown by more dark green colors near the right-bottom corner of each heatmap.
 For example, when conducting Python-to-Rust translation and JavaScript-to-Go translation on medium-level tasks, the improvement on Qwen32B is 13.64% and 8.96%, respectively; in comparison, it is only 5.54% and 4.05% in the opposite direction, respectively.
@@
 The pseudocode-based translation enables LLMs to generate Rust code in this manner, which may better align with the training corpus.
-Also, translating from Rust (top row in heatmaps) gains noticeable improvement in many trials, suggesting that pseudocode may help the understanding of original programs written in Rust.
-The results demonstrate the helpfulness of pseudocode in assisting code translation involving a low-resource PL.
+Also, translating from Rust (top row in heatmaps) gains noticeable improvement in many trials, suggesting that pseudocode also helps the understanding of original programs written in Rust.
+The results demonstrate the helpfulness of pseudocode in assisting code translation involving a low-training-resource PL.
 Message 3: Pseudocode-based translation generally improves code translation accuracy across most PL pairs.
-Besides, the improvement is more significant when translating from flexible and lightweight PLs to strict and complex PLs, and when translation involves the low-resource Rust.
+In addition, improvement is more significant when translating from flexible and lightweight PLs to strict and complex PLs, and when dealing with the low-training-resource Rust.
 Practitioners are highly advised to consider pseudocode-based translation in these scenarios.
@@
 Since no universally effective intermediate PL has been identified (Macedo et al., 2025), we implement the baseline for each studied PL and compare our $\mathcal{D}\&\mathcal{PC}$ strategy against each $\mathcal{D}\&\mathcal{PL}_{L_{i}}$, where $L_{i}\in\left\{\text{Python},\text{JavaScript},\text{C++},\text{Java},\text{Go},\text{Rust}\right\}$, with pass@10 as the metric.
-(a) vs. $\mathcal{D}\&\mathcal{PL}_{\text{Python}}$ (Qwen-32B)
-(b) vs. $\mathcal{D}\&\mathcal{PL}_{\text{JavaScript}}$ (Qwen-32B)
-(c) vs. $\mathcal{D}\&\mathcal{PL}_{\text{C++}}$ (Qwen-32B)
-(d) vs. $\mathcal{D}\&\mathcal{PL}_{\text{Java}}$ (Qwen-32B)
-(e) vs. $\mathcal{D}\&\mathcal{PL}_{\text{Go}}$ (Qwen-32B)
-(f) vs. $\mathcal{D}\&\mathcal{PL}_{\text{Rust}}$ (Qwen-32B)
-Figure 5. Relative Improvement in Pass@10 Rates of Qwen-32B with Hybrid Strategy based on Pseudocode Compared to an Intermediate PL. (The greener, the more improvement. Red value means accuracy drop.)
+(a) vs. $\mathcal{D}\&\mathcal{PL}_{\text{Python}}$ (Qwen32B)
+(b) vs. $\mathcal{D}\&\mathcal{PL}_{\text{JavaScript}}$ (Qwen32B)
+(c) vs. $\mathcal{D}\&\mathcal{PL}_{\text{C++}}$ (Qwen32B)
+(d) vs. $\mathcal{D}\&\mathcal{PL}_{\text{Java}}$ (Qwen32B)
+(e) vs. $\mathcal{D}\&\mathcal{PL}_{\text{Go}}$ (Qwen32B)
+(f) vs. $\mathcal{D}\&\mathcal{PL}_{\text{Rust}}$ (Qwen32B)
+Figure 6.
+Relative Improvement in Pass@10 Rates of Qwen32B Using Hybrid Strategy Based on Pseudocode Compared to Using Hybrid Strategy Based on an Intermediate PL.
+(The greener, the more improvement.
+Red value means performance drop.)
 Figure 6 presents the comparison results of pseudocode-based translation ($\mathcal{D}\&\mathcal{PC}$) versus intermediate PL-based translation ($\mathcal{D}\&\mathcal{PL}_{L_{i}}$) on Qwen32B.
-The results show that pseudocode-based translations achieve higher code translation accuracy than intermediate PL-based translations in most cases.
+The results show that pseudocode-based translations achieve higher code translation accuracy than intermediate PL-based translations in most situations.
 Specifically, among all 120 comparisons ($5$ source PLs $\times$ 44 target PLs $\times$ 66 transitive PLs), pseudocode-based translation wins in 109 comparisons (green cells), whereas it loses in only 11 comparisons (red cells) with marginal differences (less than 1%).
 Although certain intermediate PLs are more helpful than pseudocode for some source-target PL pairs (e.g., using JavaScript as the intermediate PL for Rust-to-Python, Go-to-C++, Go-to-Java, and C++-to-Java translation, Figure 6(b)), the pseudocode-based translations yield a generally more pronounced outperformance.
-The observed trends and conclusions are consistent across the other experimental LLMs (i.e., Qwen7B, Phi-4, GPT-4o-mini, and DSCoder).
+The observed trends and conclusions are consistent across the other studied LLMs (i.e., Qwen7B, Phi-4, GPT-4o-mini, and DSCoder).
 To avoid redundancy, we omit their detailed results here.
-These results are available in our artifact (Artifact of this paper, [n.
-d.]).
+These results are available in our artifact (Artifact of this paper, ).
 Through case studies, we found that the advantage of pseudocode over PLs is its role as a general abstraction to bridge differences across diverse source-target PL pairs.
@@
 An inappropriate choice of intermediate PL may even increase the gap between the source and target PLs and introduce extra burdens in code translation, hindering the translation accuracy.
-Besides, an intermediate PL may also bring LLMs new burdens to satisfy its syntax and semantics, complicating the translation process.
-In contrast, flexible and descriptive pseudocode describes code intent and logic, offering a PL-agnostic abstraction that guides LLMs to focus on semantics instead of detailed PL-specific implementations, thereby offering better effectiveness in general.
-We also examined cases where pseudocode underperformed PLs in transitive translation and found that inferiority mainly stemmed from the missing essential details in the pseudocode (which will be discussed in Section 4.5, Limit-2).
-In comparison, programs in intermediate PLs often preserve the complete code semantics, facilitating the generation of a semantically consistent translation result.
+In addition, an intermediate PL may also bring LLMs new burdens to satisfy its syntax and semantics, complicating the translation process.
+In contrast, flexible and descriptive pseudocode presents code intent and logic, offering a PL-agnostic abstraction that guides LLMs to focus on semantics instead of detailed PL-specific implementations, thereby offering better effectiveness in general.
+We also examined cases where pseudocode underperformed PLs in transitive translation and found that inferiority mainly stemmed from missing essential details in the pseudocode (which will be discussed in Section 4.5.2, Limit-2).
+In comparison, programs in effective intermediate PLs often preserve the complete code semantics, facilitating the generation of a semantically preserving translation result.
 These findings highlight the advantages of pseudocode as a general effective intermediary, as well as the need for enhancing its quality to fully harness its effectiveness.
 Message 4: Pseudocode serves as a more general and effective intermediary than a specific PL in transitive code translation for most studied PL pairs.
-This is because pseudocode provides a PL-agnostic abstraction of code semantics that generally bridges differences across diverse source-target PL pairs, as an effective transitive PL acts for their fitting PL pairs.
+This is because pseudocode provides a PL-agnostic abstraction of code semantics that generally bridges differences across diverse source-target PL pairs, as an effective transitive PL does for suitable PL pairs.
+In development, practitioners can prioritize pseudocode as a universal bridge for translation across various PLs.
 4.4. RQ4: Effectiveness of Code Translation based on High-Quality Pseudocode
@@
 Such low-quality pseudocode can mislead code translation.
-This RQ therefore examines whether better performance can be achieved when high-quality pseudocode is provided and identifies the bottlenecks in pseudocode-based code translation.
-Specifically, Wu et al.
-(2025) demonstrate that DeepSeek-R1 can help annotate human-written-like pseudocode with high accuracy and naturalness for concrete programs.
-Thus, in this RQ, we take DeepSeek-R1 as a high-quality pseudocode generator [^1] To avoid potential data contamination of DeepSeek-R1, whose knowledge cutoff date is unknown, we only adopt it to prepare high-quality pseudocode following Wu et al.
+This RQ therefore examines whether better performance can be achieved when high-quality pseudocode is available, as well as identifies the bottlenecks in pseudocode-based code translation.
+Specifically, DeepSeek-R1 (DeepSeek-AI, 2025) is found capable of helping annotate human-written-like pseudocode with high accuracy and naturalness for concrete programs (Wu et al., 2025).
+Therefore, in this RQ, we take DeepSeek-R1 as a high-quality pseudocode generator [^2] To avoid potential data contamination of DeepSeek-R1, whose knowledge cutoff date is unknown, we only adopt it to prepare high-quality pseudocode following Wu et al.
 (2025).
-We do not include it as a translator LLM to study..
-Then, we evaluate the translation accuracy of the studied LLMs based on DeepSeek-R1-generated high-quality pseudocode.
+We do not include it as a translator LLM to study.
+and evaluate the translation accuracy of the studied LLMs based on DeepSeek-R1-generated high-quality pseudocode.
 The investigation results help indicate the potential effectiveness of pseudocode-based translation when higher-quality pseudocode is available (e.g., written by humans or generated by enhanced pseudocode generation methods).
-Table 3. Pass@10 of Code Translation Based on Pseudocode Generated by Translator LLM (Self-Gen) and DeepSeek-R1 (R1-Generated).
+Table 4. Pass@10 of Round-Trip Code Regeneration Based on Pseudocode Generated by Translator LLM (Self-Gen) and DeepSeek-R1 (R1-Gen).
+LLM Easy Tasks Medium Tasks Hard Tasks
+LLM-Gen R1-Gen LLM-Gen R1-Gen LLM-Gen R1-Gen
+Qwen32B 0.9516 0.9670 (+1.62%) 0.8976 0.9042 (+0.74%) 0.7352 0.8050 (+9.49%)
+Phi-4 0.8889 0.9425 (+6.03%) 0.8010 0.8365 (+4.43%) 0.6738 0.6773 (+0.52%)
+GPT-4o-mini 0.8865 0.9583 (+8.10%) 0.7583 0.8646 (+14.02%) 0.5532 0.6986 (+26.28%)
+Qwen7B 0.8008 0.9267 (+15.72%) 0.6413 0.8156 (+27.18%) 0.3865 0.5567 (+44.04%)
+DS-Coder 0.8774 0.9353 (+6.60%) 0.7833 0.8188 (+4.53%) 0.5875 0.6525 (+11.06%)
+(+x.xx%): performance improvement of a strategy based on R1-generated pseudocode compared to translator-LLM-generated pseudocode.
+Quality Check of R1-Generated Pseudocode: Before comparing the code translation accuracy based on different pseudocode, we verify the assumed higher quality of pseudocode generated by DeepSeek-R1 over the studied LLMs using round-trip validation, an effective method to validating quality of semantic representations like code summaries (Allamanis et al., 2024), natural language translations (Aiken and Park, 2010), and pseudocode (Wu et al., 2025).
+Specifically, we use each studied LLM to regenerate the original program based on the pseudocode generated by itself and DeepSeek-R1, respectively, following the prompt in Figure 3(b) with the original PL.
+The regeneration pass rates can indicate the pseudocode quality, since higher-quality pseudocode should better describe code intent and logic of original programs, thus facilitating regeneration.
+As shown in Table 4, the pass rates based on R1-generated pseudocode are consistently higher than translator-generated pseudocode, suggesting the higher quality of R1-generated pseudocode in describing the code intent and logic of original programs.
+Improvement with High-quality Pseudocode.
+Table 5 compares the pass@10 rates of different translation strategies based on high-quality pseudocode generated by DeepSeek-R1, and pseudocode generated by the studied LLMs themselves (i.e., the default setup studied in previous RQs), respectively.
+The results demonstrate that the code translation accuracy consistently improves across LLMs and strategies after switching to DeepSeek-R1-generated pseudocode.
+Among all strategies, the pure pseudocode-based strategy ($\mathcal{P}$) benefited the most, with average improvements of 8.16%, 13.45%, and 22.15% on three difficulty levels, respectively, mitigating the underperformance relative to direct translation ($\mathcal{D}$) observed in RQ1.
+These results indicate pseudocode quality as a critical bottleneck for pseudocode-based translation.
+In addition, we observed that hybrid strategies ($\mathcal{D}\&\mathcal{P}$ and $\mathcal{D}\&\mathcal{PC}$) continuously outperformed single-approach strategies ($\mathcal{D}$, $\mathcal{P}$, $\mathcal{PC}$), indicating that combining direct and pseudocode-based translation remains beneficial even with higher-quality pseudocode.
+With DeepSeek-R1-generated high-quality pseudocode, the best pass@10 scores of all LLMs with optimal strategy improved to 0.9646-0.9835, 0.8861–0.9512, and 0.6747–0.8286 across three difficulty levels, demonstrating the promising potential of pseudocode-based code translation.
+Table 5. Pass@10 of Code Translation Based on Pseudocode Generated by Translator LLM (Self-Gen) and DeepSeek-R1 (R1-Gen).
 LLM Strategy Easy Tasks Medium Tasks Hard Tasks
-Self-Gen R1-Generated Self-Gen R1-Generated Self-Gen R1-Generated
+Self-Gen R1-Gen Self-Gen R1-Gen Self-Gen R1-Gen
 Qwen32B $\mathcal{D}$ 0.9480 - - 0.8772 - - 0.7059 - -
@@
 $\mathcal{D}\&\mathcal{P}$ 0.9646 0.9781 (+1.40%) 0.8947 0.9192 (+2.74%) 0.7083 0.7563 (+6.78%)
-$\mathcal{D}\&\mathcal{PC}$ 0.9610 0.9759 (+1.55%) 0.8903 0.9127 (+2.52%) 0.7038 0.7385 (+4.93%)
+$\mathcal{D}\&\mathcal{PC}$ 0.9610 0.9759 (+1.55%) 0.8903 0.9127 (+2.52%) 0.7040 0.7385 (+4.90%)
 GPT-4o -mini $\mathcal{D}$ 0.9279 - - 0.8552 - - 0.6667 - -
@@
 $\mathcal{D}\&\mathcal{PC}$ 0.9537 0.9726 (+1.98%) 0.8581 0.8954 (+4.35%) 0.6459 0.7073 (+9.51%)
-Average $\mathcal{D}$ 0.9214 - - 0.8288 - - 0.6199 - -
+(Average) $\mathcal{D}$ 0.9214 - - 0.8288 - - 0.6199 - -
 $\mathcal{P}$ 0.8634 0.9338 (+8.16%) 0.7392 0.8386 (+13.45%) 0.5191 0.6341 (+22.15%)
@@
 $\mathcal{D}\&\mathcal{P}$ 0.9579 0.9745 (+1.73%) 0.8874 0.9158 (+3.20%) 0.6945 0.7465 (+7.49%)
-$\mathcal{D}\&\mathcal{PC}$ 0.9592 0.9750 (+1.64%) 0.8905 0.9149 (+2.74%) 0.7051 0.7481 (+6.10%)
-(+xx%): performance improvement of a strategy based on R1-generated pseucode compared to LLM-translator-generated pseudocode.
-Improvement with High-quality Pseudocode.
-Table 3 compares the pass@10 rates of different translation strategies based on high-quality pseudocode generated by DeepSeek-R1 and pseudocode generated by the studied LLMs themselves (i.e., the default setup studied in previous RQs), respectively.
-The results show that the code translation accuracy consistently improves across LLMs and strategies when using R1-generated pseudocode.
-Among all strategies, the pure pseudocode-based strategy ($\mathcal{P}$) benefits the most, with average improvements of 8.16%, 13.45%, and 22.15% on three difficulty levels, respectively, mitigating the underperformance relative to direct translation ($\mathcal{D}$) observed in RQ1.
-These results indicate pseudocode quality as a critical bottleneck for pseudocode-based translation.
-In addition, we observed that hybrid strategies ($\mathcal{D}\&\mathcal{P}$ and $\mathcal{D}\&\mathcal{PC}$) continuously outperform single-approach strategies ($\mathcal{D}$, $\mathcal{P}$, $\mathcal{PC}$), indicating that combining direct and pseudocode-based translation remains beneficial even with higher-quality pseudocode.
-With R1-generated high-quality pseudocode, the best pass@10 scores of all LLMs with optimal strategy improved to 0.9646-0.9835, 0.8861–0.9512, and 0.6747–0.8286 across the three difficulty levels, demonstrating the promising potential of pseudocode-based code translation.
-Message 5: The quality of pseudocode hinders accurate pseudocode-based translation of the studied LLMs.
+$\mathcal{D}\&\mathcal{PC}$ 0.9592 0.9750 (+1.64%) 0.8905 0.9149 (+2.74%) 0.7052 0.7481 (+6.09%)
+(+x.xx%): performance improvement of a strategy based on R1-generated pseudocode compared to translator-LLM-generated pseudocode.
+Message 5: The quality of pseudocode hinders accurate pseudocode-based code translation of the studied LLMs.
 Higher-quality pseudocode can consistently improve the code translation accuracy of studied LLMs across translation strategies and task difficulties.
+Practitioners may also consider human-in-the-loop to prepare high-quality pseudocode or revise the generated pseudocode, which will help guide LLMs to generate more accurate translation results through pseudocode-based code translation.
 LLM-wise Comparison.
-The performance comparison among LLMs further demonstrates the remaining limitations of LLMs in both code understanding and generation during pseudocode-based code translation.
-Specifically, we observed that the performance improvement brought by R1-generated pseudocode is more significant for weaker LLMs.
+The performance comparison among LLMs further demonstrates the limitations of LLMs in both code understanding and generation during pseudocode-based code translation.
+Specifically, we observed that the performance improvement brought by DeepSeek-R1-generated pseudocode is more significant for weaker LLMs.
 For example, the improvement of all strategies on Qwen7B ranges from 2.31% to 64.03%, while it is only 0.61% to 8.98% on Qwen32B, indicating that higher-quality pseudocode is more beneficial for weaker LLMs.
-This demonstrates the weakness of weaker LLMs in understanding and summarizing the code intent and logic of original programs.
-In addition, with R1-generated pseudocode, the performance gap between weaker LLMs and more powerful LLMs remains.
+This demonstrates the weakness of weaker LLMs in understanding and interpreting the code intent and logic of original programs.
+In addition, with DeepSeek-R1-generated pseudocode, the performance gap between weaker LLMs and more powerful LLMs remains.
 For example, with R1-generated pseudocode, the pass@10 rates of Qwen7B and DSCoder with $\mathcal{D}\&\mathcal{PC}$ are only 0.6747 and 0.7073 on hard tasks, respectively, which are much lower than 0.8286 of Qwen32B.
 The results indicate that even given the same code intent and logic (i.e., pseudocode), the weaker LLMs are still less capable in code translation than the more powerful LLMs, indicating that the code implementation capability also hinders the effectiveness of pseudocode-based code translation on weaker LLMs.
-Message 6: Both the code understanding capability (indicated by the quality of the generated pseudocode) and the code implementation capability of the studied LLMs hinder their performance in code translations.
+Message 6: Both code understanding capability (indicated by the quality of the interpreted pseudocode) and code implementation capability (based on pseudocode) of the studied LLMs hinder their performance in code translation tasks.
+Further advancements in generating high-quality pseudocode from programs and implementing accurate code from pseudocode are necessary to fully unleash the potential of pseudocode-based code translation.
 4.5. Discussion: Advantages and Limitations of Pseudocode-Based Translation
 After understanding the effectiveness of pseudocode-based code translation from the quantitative results in four RQs, we further conduct case studies to attribute its successes and failures to learn concrete insights for the effective adoption of this approach.
-Based on these findings, we also suggest several research directions to mitigate the limitations and harness the potential.
-Key advantages of pseudocode-based code translation.
-We identified three major advantages of emulating explicit semantic translation via pseudocode-based translation to complement direct translation.
-One of them facilitates the understanding of the original programs and reduces the burden of one step, and the other two help LLMs handle differences across PLs.
-Practitioners are encouraged to leverage the advantages of pseudocode-based translation for original programs or PL pair fitting them.
+Specifically, two authors with over five years of programming experience analyzed the results of 200 randomly sampled translation tasks (i.e., 100 tasks where direct translation failed while pseudocode-based translation succeeded, and another 100 tasks where direct translation succeeded while pseudocode-based translation failed [^3] To study reasons of clear successes/failures rather than randomness, we only analyzed the tasks where over five in ten attempts of a strategy passed all tests (successful tasks) and where none attempt passed all tests (failed tasks).), to identify the advantages/limitations of pseudocode for successful/failed translations compared with direct translation.
+These tasks are sampled across all the studied LLMs and PL pairs, representing a 95% confidence level with a 10% margin of error.
+The participants first discussed potential categorizations and criteria of advantages and limitations via a pilot study on 40 sampled tasks.
+Then, they independently annotated each task (with substantial agreement indicated by Cohen’s Kappa coefficients (Cohen, 1960) over 0.78), and next discussed to reach a consensus on all the annotations.
+Each task can be annotated with multiple advantages/limitations of pseudocode.
+In addition, based on all the findings in this work, we further discuss several research directions to mitigate the limitations and harness the potential for pseudocode-based code translation.
+4.5.1. Key Advantages of Pseudocode-based Code Translation.
+We identified three major advantages of pseudocode-based translation over direct translation.
+One of them facilitates the understanding of the original programs and reduces the burden of an integrated step, and the other two help LLMs handle differences across PLs.
+They align with the limitations of direct translation identified in prior studies (Pan et al., 2024; Jiao et al., 2023; Macedo et al., 2025).
+Practitioners are encouraged to leverage these advantages to translate the original programs or across PL pairs fitting them.
   1. Pro-1
-    Pseudocode helps disentangle code understanding and implementation in code translation into two steps, formulating easier subproblems for LLMs to solve over complicated programs.
-Some programs are compact, implementing multiple operations in a single dense statement by chaining several steps recursively or sequentially (Keshk and Dyer, 2023).
+    Pseudocode helps decouple code understanding and implementation in code translation, formulating easier subproblems for LLMs to solve when translating complicated programs.
+Some programs are compact, implementing multiple operations in a single dense statement by chaining several steps recursively or sequentially.
 Such programs may result from PL idioms or developers’ programming style (Keshk and Dyer, 2023).
-We observed that LLMs struggle to disentangle the functionality of such complicated programs and re-implement the functionality in the target PL via one-step direct translation.
+We observed that LLMs struggle to understand the tangled functionality of such complicated programs and re-implement the functionality in the target PL via one-step direct translation.
 LLMs often miss or mess up steps in such cases.
 In comparison, a pseudocode-based transitive translation splits the code understanding of the complicated program and the implementation of the code intent and logic into two steps, reducing the burden on LLMs and promoting accuracy in each step.
+This advantage was observed in 34% of analyzed cases where pseudocode-based translation succeeded while direct translation failed.
 Figure 7 illustrates an example of this case.
-Specifically, the original program includes only one Python statement that recursively conducts four operations (LABEL:lst:_rq1-type3-py2js-python).
+Specifically, the original Python program includes only one statement that recursively conducts four operations (LABEL:lst:_rq1-type3-py2js-python).
 When translating this compact program via the one-step direct translation, even the powerful Qwen32B missed the row-sorting step in the output JavaScript program (LABEL:lst:_rq1-type3-py2js-direct).
-In comparison, when asked to describe the code intent and logic using pseudocode, Qwen32B was able to interpret the complicated logic in the compact Python program in pseudocode (LABEL:lst:_rq1-type3-py2js-pseudocode).
-The pseudocode then promotes the successful implementation of the operation JavaScript (LABEL:lst:_rq1-type3-py2js-viapseudocode).
-Similar symptoms are also observed on Phi-4 and Qwen7B.
+In comparison, when asked to describe the code intent and logic using pseudocode, Qwen32B was able to interpret the complicated logic of the compact Python code using pseudocode (LABEL:lst:_rq1-type3-py2js-pseudocode).
+The pseudocode then promotes the successful implementation of the operation in JavaScript (LABEL:lst:_rq1-type3-py2js-viapseudocode).
+Phi-4 and Qwen7B also suffered from similar issues when conducting direct translation and succeeded in pseudocode-based translation.
 def matrixSum(self, nums: List[List[int]]) -> int:
@@
 # val extraction, and sum in a single expression.
-Listing 1: Source Python Program to Translate
+Listing 1: Original Python Program to Translate
 function matrixSum(nums)
@@
 return sum of max_values
-Listing 2: Pseudocode of Source Python Program
+Listing 2: Pseudocode of Original Python Program
 var matrixSum = function(nums) {
@@
 ).reduce((sum, maxVal) => sum + maxVal, 0);
-Listing 3: Translation with Direct
+Listing 3: Direct Translation Result
 var matrixSum = function(nums) {
@@
 const transposedMatrix = sortedRows[0].map(...
-Listing 4: Translation with viaPseudocode (w/ code)
-Figure 6. An Example of Python-to-JavaScript Translation Involving Compact Original Programs (Qwen32B)3
+Listing 4: Pseudocode-based Translation Result
+Figure 7. An Example of Python-to-JavaScript Translation Involving Compact Original Programs (Qwen32B)5
   1. Pro-2
     Pseudocode abstracts the PL-specific features (e.g., variable scoping, closure capture mode, and syntax sugar), avoids LLMs in following the incompatible practices in original programs.
-We observed that LLM-based direct translation may overlook the differences in the features of source and target PLs, resulting in incompatible implementations violating the mechanism of the target PL.
-Meanwhile, pseudocode describes the code’s intent and logic, hiding PL-specific implementation.
-When generating target code from the pseudocode, LLMs often properly implement the intended function in a manner fitting the mechanism of the target PL, and are less likely to be distracted by the incompatible patterns of source PL in the original program.
-Footnote 2 shows an example of this case with a Go-to-Rust translation.
-Specifically, when translating a Go program with an inner function (LABEL:lst:_rq1-type1-go2rust-go) into Rust via direct translation, Qwen7B mechanically followed the Go program to declare a common inner function in Rust without realizing the difference in default closure capture modes in Go and Rust and forgetting to indicate a mutable closure capture mode (LABEL:lst:_rq1-type1-go2rust-direct).
-Meanwhile, based on pseudocode that mainly indicates the code intent and logic (LABEL:lst:_rq1-type1-go2rust-pseudocode), Qwen7B correctly implemented the functionality in a proper manner that fits Rust’s memory management (LABEL:lst:_rq1-type1-go2rust-viapseudocode).
-Interestingly, all five studied LLMs failed with the same fault in direct translation while succeeding with the help of pseudocode.
+We observed that LLM-based direct translation may overlook the differences in the features of source and target PLs, resulting in incompatible implementations in the translation result that violate the mechanism of the target PL.
+Meanwhile, pseudocode describes the original program’s intent and logic, hiding PL-specific implementation.
+When generating target code from pseudocode, LLMs often properly implement the intended function in a manner fitting the mechanism of the target PL, and are less likely to be distracted by the incompatible patterns of source PL in original programs.
+This advantage was observed in 56% of analyzed cases where pseudocode-based translation outperformed direct translation.
+Footnote 4 shows an example of this case with a Go-to-Rust translation.
+Specifically, when translating a Go program with an inner function (LABEL:lst:_rq1-type1-go2rust-go) into Rust via direct translation, Qwen7B mechanically followed the Go program to declare a common inner function in Rust without realizing the difference in default closure capture modes in Go and Rust and failing to indicate a mutable closure capture mode in Rust (LABEL:lst:_rq1-type1-go2rust-direct).
+Meanwhile, based on pseudocode that mainly indicates the code intent and logic (LABEL:lst:_rq1-type1-go2rust-pseudocode), Qwen7B correctly implemented the functionality in a proper manner that fits the memory management of Rust (LABEL:lst:_rq1-type1-go2rust-viapseudocode).
+Notably, all five studied LLMs failed with the same fault in direct translation while succeeding with the help of pseudocode.
 This also echoes the helpfulness of pseudocode-based code translation in translating from flexible PLs to more rigid PLs observed in RQ2 (Section 4.2).
@@
 if r == ’W’ { res++; } //capture by default
-Listing 5: Source Go Program to Translate
+Listing 5: Original Go Program to Translate
 function canMakeSquare(grid):
 count = count W in the 2x2 block
-Listing 6: Pseudocode of Source Go Program
+Listing 6: Pseudocode of Original Go Program
 pub fn can_make_square(grid: Vec<Vec<char>>) -> bool {
@@
 res += 1;//, thus illegal to mutate res
-Listing 7: Translation with Direct
+Listing 7: Direct Translation Result
 pub fn can_make_square(grid: Vec<Vec<char>>) -> bool {
@@
 count += 1;
-Listing 8: Translation with viaPseudocode (w/ code)
+Listing 8: Pseudocode-based Translation Result
 Figure 8.
-An Example of Go-to-Rust Translation Involving Unique Closure Capturing in Go and Rust (Qwen7B)333“…” represents code less related to the illustration and omitted for saving space.
+An Example of Go-to-Rust Translation Involving Unique Closure Capturing in Go and Rust (Qwen7B)555“…” represents code less related to the illustration and omitted for saving space.
 The comment is not generated by LLMs but annotated by authors for case explanation.
@@
 We observed that LLMs could hallucinate incorrect implementations that fail to realize the functionality of the PL-specific APIs in original programs during direct translation.
-LLMs struggle to identify or implement a semantically equivalent API counterpart in the target PL based on only the original program, and may even hallucinate non-existing APIs.
+LLMs struggle to identify or implement a semantically equivalent API counterpart in the target PL based on only the original program, and may even hallucinate non-existent APIs.
 Meanwhile, pseudocode often describes the functionality implemented with PL-specific APIs in natural language, guiding LLMs to figure out appropriate APIs and API usages in the target PL as they work out a code generation task without being misled by concrete API usages in original programs.
+This advantage was observed in 33% of analyzed cases where pseudocode-based translation outperformed direct translation.
 Figure 9 illustrates an example of this case.
@@
 However, the generated bit operations do not align with the functionality of “bit_count”.
-Similarly, GPT-4o-mini hallucinated incorrect bit operations, and DSCoder even hallucinated a non-existing API.
+Similarly, GPT-4o-mini hallucinated incorrect bit operations, and DSCoder even hallucinated a non-existent API.
 Instead, the pseudocode generated by Phi-4 clarified the functionality implemented by “bit_count” in Python (LABEL:lst:_rq1-type2-py2go-pseudocode), which guided Phi-4 to generate a correct subroutine “countBits” in Go to realize the functionality of “bit_count” (LABEL:lst:_rq1-type2-py2go-viapseudocode).
@@
 return ascii_lowercase[(k-1).bit_count()]
-Listing 9: Source Python Program to Translate
+Listing 9: Original Python Program to Translate
 function kthCharacter(k)
@@
 return character at index from string_of_lowercase_letters
-Listing 10: Pseudocode of Source Python Program
+Listing 10: Pseudocode of Original Python Program
 func kthCharacter(k int) byte {
@@
 // DSCoder:
-// index := (k - 1).bit_count() //non-existing API
-Listing 11: Translation with Direct
+// index := (k - 1).bit_count() //non-existent API
+Listing 11: Direct Translation Result
 func kthCharacter(k int) byte {
@@
 func countBits(n int) int {
-count := 0
-for n > 0 {
+count := 0 // Counting bits in the integer as
+for n > 0 { // the Python API bit_count does.
 count += n & 1
@@
 return count
-Listing 12: Translation with viaPseudocode (w/ code)
-Figure 8. An Example of Python-to-Go Translation Involving Non-Existing Equivalent APIs in Go (Phi-4)3
-Typical limitations of pseudocode-based code translation.
-We also observed three major limitations of pseudocode that cause errors in semantic code translation results.
+Listing 12: Pseudocode-based Translation Result
+Figure 9. An Example of Python-to-Go Translation Involving Non-Existent Equivalent APIs in Go (Phi-4)5
+4.5.2. Typical Limitations of Pseudocode-based Code Translation.
+We also observed three major limitations that cause errors in pseudocode-based code translation due to issues in the generated pseudocode.
 They stem from LLMs’ incorrect understanding of original programs, the semantic loss in information transmission, and the ambiguity of natural language.
-Enhancements to generate more accurate pseudocode are needed to mitigate these issues and harness the advantages of pseudocode-based code translation.
+Enhancements to generate more accurate pseudocode are needed to mitigate these limitations and harness the advantages of pseudocode-based code translation.
   1. Limit-1
@@
 For example, translating a straightforward Python expression “sum = sum+nums[i] if nums[i]<=sum else nums[i]” into Java may not necessitate the understanding of the code functionality.
-An incorrect intent conversely misled translation: Qwen32B hallucinated a code intent of “sum = max(sum, sum + nums[i])” for this expression, misleading the generation of the Java program.
+An incorrect intent conversely misled translation: Qwen32B hallucinated a code intent of “sum = max(sum, sum + nums[i])” for this expression in pseudocode, misleading the generation of the Java program.
 The issue happened with both $\mathcal{P}$ and $\mathcal{PC}$ strategies, while direct translation ($\mathcal{D}$) leads to a correct translation.
+This limitation was observed in 66% of analyzed cases where pseudocode-based translation failed while direct translation succeeded.
   1. Limit-2
@@
 Another typical missing information for reproducing code semantics is the logic of complicated algorithms.
-From a concise description without detailed implementations, LLMs may not accurately reproduce all algorithms.
+From a brief description of logic without detailed implementations, LLMs may not accurately reproduce all algorithms.
 The concise description may also miss customized operations in algorithms, leaving the generated code degenerate into a general implementation lacking the expected features.
+This limitation was observed in 27% of analyzed cases where direct translation outperformed pseudocode-based translation.
   1. Limit-3
@@
 Although natural language can depict PL-agnostic code intent and logic, its nature of ambiguity may introduce noise and mislead the code generation step.
-A typical issue observed in the failure cases is the ambiguity of “to” and “downto” in “for” loops.
-Such expressions cannot clarify the inclusion of the border value in the loop.
-The inclusion of the original program (i.e., using $\mathcal{PC}$ strategy) cannot always mitigate the disturbance caused by the ambiguity as well.
+A typical ambiguity is about the descriptions of loop boundaries using “to”/“downto”, which cannot clarify the inclusion of border values.
+For example, Qwen32B generated a pseudocode with “for i from 0 to length of nums” for the original Rust loop “for i in 0..nums.len()”, where the upper border value is excluded.
+When implementing the pseudocode in Go, Qwen32B generated a loop “for i:=0; i<=len(nums) …” with the upper border value included, resulting in a semantic drift.
+Another common issue is the unclear precedence among arithmetic operations described by natural language descriptions.
+For example, Qwen32B generated a pseudocode using “a + integer division of (k - 1) by a - 1” to describe the original Java expression “a+(k-1)/a-1”.
+The unclear precedence of addition and subtraction misled Qwen32B to implement “a+Math.floor((k-1)/(a-1))” in JavaScript, which drifts from original semantics.
+Including original programs (i.e., using $\mathcal{PC}$ strategy) cannot mitigate the disturbance caused by the ambiguity as well.
 In comparison, a direct translation from the original program does not suffer from such ambiguity-induced issues.
-Research Opportunities.
+This limitation was observed in 17% of analyzed cases where direct translation outperformed pseudocode-based translation.
+4.5.3. Research Opportunities.
 To leverage the advantages of pseudocode-based code translation observed in the previous three RQs and its potential revealed in RQ4, we suggest four directions for future enhancement based on our findings.
@@
 Effective validation strategies may help identify inconsistencies and semantic losses.
-For example, round-trip validation (Aiken and Park, 2010; Zhou and Sun, 2018) may identify buggy pseudocode by translating the generation result back to the source PL, and LLMs may infer the intended behavior based on code (Xu et al., 2025) and compare pseudocode to it to debug pseudocode.
-Incorporating such validation loops would help identify and mitigate failures stemming from low-quality pseudocode suffering from the three limitations.
+For example, round-trip validation (Aiken and Park, 2010; Zhou and Sun, 2018) may help identify buggy pseudocode by translating the generation result back to the source PL for comparison, and LLMs may infer the intended behavior based on code (Xu et al., 2025) and compare pseudocode to it to debug pseudocode.
+Incorporating such validation steps in the loop would help identify and mitigate failures stemming from low-quality pseudocode suffering from the three limitations.
   1. Chance-3
@@
 Such integration may help integrate the fine-grained advantages of both approaches.
+  1. Chance-5
+    Enhancing code translation via reasoning.
+Several advanced LLMs (e.g., DeepSeek-R1) have a built-in mechanism to generate reasoning traces in the response process, which proves helpful in guiding LLMs to solve complicated tasks like mathematical and programming problems (Guo et al., 2025).
+However, it remains underexplored how to effectively leverage reasoning traces for code translation.
+There can be several potential directions.
+One idea is to directly use reasoning traces as an intermediate representation to guide code translation, similar to pseudocode.
+However, reasoning traces are often more verbose and less structured than pseudocode, which may introduce unique challenges.
+For example, we observed serious overthinking issues when prompting DeepSeek-R1 Distilled Qwen7B to generate reasoning traces before code translation, leading to a decrease in translation accuracy compared to direct and pseudocode-based translation and an increase in computation cost, like issues revealed by existing studies (Chen et al., 2025b).
+In addition, the adoption of reasoning during pseudocode generation is also worth exploring to help LLMs better understand and depict original programs.
+In this study, we have employed chain-of-thought prompting (Wei et al., 2022) to trigger reasoning traces of the five LLMs studied, aiming to guide LLMs to explicitly put down the abstraction and analysis steps that may help generate correct and concise pseudocode.
+Nevertheless, more systematic designs to effectively integrate reasoning into code translation are worth further exploration.
 5. Threats to Validity
-Table 4. Pass@10 Rates on All 323 Tasks and 114 Newer Tasks After Cut-off Date (2024-06-01)
-Strategy LLM All 323 Tasks
-114 Newer Tasks
-LLM All 323 Tasks
-114 Newer Tasks
-$\mathcal{D}$(Direct Translation) Qwen 32B 0.8777 0.8604 Qwen 7B 0.7934 0.7623
-$\mathcal{P}$(viaPseudocode) 0.8703 (-0.84%) 0.8424 (-2.09%) 0.7446 (-21.43%) 0.5320 (-30.21%)
-$\mathcal{PC}$(viaPseudocode w/ prog.) 0.9023 (+2.80%) 0.8848 (+2.84%) 0.8016 (+2.96%) 0.7435 (-2.47%)
-$\mathcal{D}\&\mathcal{P}$(Hybrid [DT&vP]) 0.9325 (+6.24%) 0.9199 (+6.92%) 0.8587 (+4.92%) 0.7956 (+4.37%)
-$\mathcal{D}\&\mathcal{PC}$(Hybrid [DT&vPw/p]) 0.9292 (+5.87%) 0.9170 (+6.58%) 0.8616 (+6.93%) 0.8151 (+6.93%)
-$\mathcal{D}$(Direct Translation) Phi-4 0.8363 0.8146 DS Coder 0.7934 0.7672
-$\mathcal{P}$(viaPseudocode) 0.7839 (-6.27%) 0.7702 (-5.45%) 0.7446 (-5.23%) 0.7185 (-6.35%)
-$\mathcal{PC}$(viaPseudocode w/ prog.) 0.8398 (+0.42%) 0.8210 (+0.79%) 0.8016 (+3.30%) 0.7792 (+1.56%)
-$\mathcal{D}\&\mathcal{P}$(Hybrid [DT&vP]) 0.8926 (+6.73%) 0.8773 (+7.70%) 0.8587 (+8.23%) 0.8330 (+8.58%)
-$\mathcal{D}\&\mathcal{PC}$(Hybrid [DT&vPw/p]) 0.8886 (+6.25%) 0.8735 (+7.23%) 0.8616 (+8.60%) 0.8359 (+8.95%)
-$\mathcal{D}$(Direct Translation) GPT-4o -mini 0.8539 0.8363 Avg 0.8309 0.8082
-$\mathcal{P}$(viaPseudocode) 0.7409 (-13.23%) 0.7155 (-14.44%) 0.7769 (-9.04%) 0.7157 (-11.44%)
-$\mathcal{PC}$(viaPseudocode w/ prog.) 0.8590 (+0.60%) 0.8520 (+1.88%) 0.8409 (+2.48%) 0.8161 (+0.98%)
-$\mathcal{D}\&\mathcal{P}$(Hybrid [DT&vP]) 0.9031 (+5.76%) 0.8857 (+5.91%) 0.8891 (+6.37%) 0.8623 (+6.70%)
-$\mathcal{D}\&\mathcal{PC}$(Hybrid [DT&vPw/p]) 0.9093 (+6.49%) 0.8962 (+7.16%) 0.8901 (+6.80%) 0.8675 (+7.35%)
 We discuss three potential threats to the validity of our study and our mitigation methods as follows.
@@
 To mitigate this threat, we followed existing work to investigate code translation across six PLs.
-These include Python, C++, Java, and JavaScript, which are widely adopted in daily development and studied in existing works (TIOBE Index, [n.
-d.]; Rozière et al., 2022; Jiao et al., 2023; Yang et al., 2024), as well as the emerging Go and Rust (Pan et al., 2024; Macedo et al., 2025).
+These include Python, C++, Java, and JavaScript, which are widely adopted in daily development and studied in existing works (TIOBE Index, ; Rozière et al., 2022; Jiao et al., 2023; Yang et al., 2024), as well as the emerging Go and Rust (Pan et al., 2024; Macedo et al., 2025).
 Our study results are expected to guide the adoption of pseudocode-based translation for these popular PLs.
-We constructed translation tasks based on LeetCode problems in LiveCodeBench (Jain et al., 2025).
-LeetCode problems are widely used in code translation studies (Wang et al., 2023a; Ahmad et al., 2023) and benchmark of LLMs’ coding ability (Hui et al., 2024; Team, 2024; Jain et al., 2025) since they provide adequate test cases to evaluate candidate solutions and have easy-to-collect multilingual solution programs.
-In addition, these tasks require common engineering and algorithm implementations useful in daily development.
+We constructed translation tasks based on various LeetCode problems in LiveCodeBench (Jain et al., 2025).
+LeetCode problems are widely used in code translation studies (Wang et al., 2023a; Ahmad et al., 2023) and benchmarking LLMs’ coding ability (Hui et al., 2024; Team, 2024; Jain et al., 2025) since they provide high-quality tests to evaluate candidate solutions and provide easy access to collect multilingual programs.
+Also, they require common engineering and algorithm implementations useful in daily development.
 Thus, we consider the study results on their solution programs should be meaningful and generalizable to practical development code.
-We used five LLMs deployable on our machine as the code translator.
-These include both general LLMs and coder LLMs, and both closed-source and open-source LLMs from four families.
-They rank a varying range (from Top-10 to 68) on the famous BigCodeBench (Zhuo et al., 2025).
-Also, the training data for these LLMs has a cutoff date with relatively little overlap with the timeline of LeetCode problems.
-Thus, we consider the results learned with these LLMs should be representable and reliable.
+We used five LLMs deployable on our machine as the code translator, including both general and coding LLMs, and both closed-source and open-source LLMs from four families.
+They rank a varying range (from Top-10 to 68 according to the record on Oct 1, 2025) on the famous BigCodeBench (Zhuo et al., 2025).
+Also, the training data for these LLMs has a cutoff date with relatively little overlap with the timeline of the LeetCode problems to construct translation tasks.
+Therefore, we consider the results learned with these LLMs should be representable and reliable.
 Data Contamination.
-A potential threat to our study result is the subject to the data contamination issue, where LLMs perform well because they have learned the ground truth during training rather than pseudocode-based translation strategies.
-Although our translation tasks based on LiveCodeBench are much newer than the subjects in the conventional code translation benchmarks as introduced in Section 3.2, there is still some overlap between the timeline of LiveCodeBench problems and the studied LLMs’ training data.
-To further verify the cause of the observed performance improvement, we re-evaluate LLMs on the solution programs of a clean subset of 114 LeetCode problems released after the latest knowledge cutoff of the studied LLMs, i.e., June 1, 2024.
-As shown in Table 4, the performance improvement of different strategies is generally consistent with that observed on all 323 tasks.
-Thus, we consider the observation based on the complete set is meaningful and provides more statistically meaningful observations with more subjects.
+A potential threat to our study result is the subject to the data contamination issue, where LLMs perform well because they have learned the ground truth during training, rather than pseudocode-based translation strategies.
+Although our translation tasks built upon LiveCodeBench are much newer than the subjects in the conventional code translation benchmarks as introduced in Section 3.2, there is still some overlap between the timeline of LiveCodeBench problems and the studied LLMs’ training data.
+To further verify the cause of the observed performance improvement, we re-evaluate LLMs on the solution programs of a clean subset of 114 problems released after the latest knowledge cutoff of the studied LLMs, i.e., June 1, 2024.
+As shown in Table 6, the performance improvement of different strategies is generally consistent with that observed on all 323 problems.
+Therefore, we consider the observation based on the complete set to be meaningful and provide more statistically meaningful observations with more subjects.
+Table 6. Pass@10 Rates on All 323 Problems and 114 Newer Problems After Cut-off Date (2024-06-01)
+Strategy LLM All 323 Probs
+114 Newer Probs
+LLM All 323 Probs
+114 Newer Probs
+$\mathcal{D}$(Direct Translation) Qwen 32B 0.8777 0.8604 Qwen 7B 0.7971 0.7623
+$\mathcal{P}$(viaPseudocode) 0.8703 (-0.84%) 0.8424 (-2.09%) 0.6193 (-22.31%) 0.5320 (-30.21%)
+$\mathcal{PC}$(viaPseudocode w/ ctx.) 0.9023 (+2.80%) 0.8848 (+2.84%) 0.7954 (-0.21%) 0.7435 (-2.47%)
+$\mathcal{D}\&\mathcal{P}$(Hybrid[DT&vP]) 0.9325 (+6.24%) 0.9199 (+6.92%) 0.8363 (+4.92%) 0.7956 (+4.37%)
+$\mathcal{D}\&\mathcal{PC}$(Hybrid[DT&vPw/c.]) 0.9292 (+5.87%) 0.9170 (+6.58%) 0.8523 (+6.93%) 0.8151 (+6.93%)
+$\mathcal{D}$(Direct Translation) Phi-4 0.8363 0.8146 DS Coder 0.7934 0.7672
+$\mathcal{P}$(viaPseudocode) 0.7839 (-6.27%) 0.7702 (-5.45%) 0.7446 (-5.23%) 0.7185 (-6.35%)
+$\mathcal{PC}$(viaPseudocode w/ ctx.) 0.8398 (+0.42%) 0.8210 (+0.79%) 0.8016 (+3.30%) 0.7792 (+1.56%)
+$\mathcal{D}\&\mathcal{P}$(Hybrid[DT&vP]) 0.8926 (+6.73%) 0.8773 (+7.70%) 0.8587 (+8.23%) 0.8330 (+8.58%)
+$\mathcal{D}\&\mathcal{PC}$(Hybrid[DT&vPw/c.]) 0.8886 (+6.25%) 0.8735 (+7.23%) 0.8616 (+8.60%) 0.8359 (+8.95%)
+$\mathcal{D}$(Direct Translation) GPT-4o -mini 0.8539 0.8363 (Avg) 0.8317 0.8082
+$\mathcal{P}$(viaPseudocode) 0.7409 (-13.23%) 0.7155 (-14.44%) 0.7518 (-9.60%) 0.7157 (-11.44%)
+$\mathcal{PC}$(viaPseudocode w/ ctx.) 0.8590 (+0.60%) 0.8520 (+1.88%) 0.8396 (+0.95%) 0.8161 (+0.98%)
+$\mathcal{D}\&\mathcal{P}$(Hybrid[DT&vP]) 0.9031 (+5.76%) 0.8857 (+5.91%) 0.8846 (+6.37%) 0.8623 (+6.70%)
+$\mathcal{D}\&\mathcal{PC}$(Hybrid[DT&vPw/c.]) 0.9093 (+6.49%) 0.8962 (+7.16%) 0.8882 (+6.80%) 0.8675 (+7.35%)
 Representativeness of Studied Translation Strategies.
-The representativeness of the studied translation strategies and prompts is another potential threat to our study, impacting the meaningfulness of our observations.
-To mitigate this threat, we studied five translation strategies, including one direct translation strategy, which is the commonly used approach in existing studies and practices, as well as two pseudocode-based translation strategies and two hybrid translation strategies.
+The representativeness of the studied translation strategies and prompts is another potential threat to our study, affecting the meaningfulness of our observations.
+To mitigate this threat, we studied five code translation strategies, including one direct translation strategy commonly used in LLM-based code translation studies and practices (Pan et al., 2024; Yang et al., 2024), as well as two pseudocode-based translation strategies and two hybrid translation strategies.
 The pseudocode-based strategies emulate the human practices of semantic translation on natural languages (Newmark, 1981) and are implemented in a transitive translation manner following (Macedo et al., 2025).
@@
 6.1. Automated Code Translation
-To facilitate development activities relying on code translation, various automated code translation methods have emerged in the past decades (Chen et al., 2025; Jiao et al., 2023).
+To facilitate development activities relying on code translation, various automated code translation methods have emerged in the past decades (Chen et al., 2025a; Jiao et al., 2023).
 Nguyen et al.
@@
 Existing automated code translation approaches mainly focus on code-to-code transformation.
-Differently, we study the effectiveness of emulating semantic translation for code via pseudocode, where LLMs first describe the semantics of original programs with pseudocode and then produce translated programs.
-We identify the complementary role of pseudocode-based translation for conventional direct translation, as well as its helpfulness in reducing task difficulties and bridging the differences between the features and APIs in varying PL pairs.
-The findings may guide practitioners to obtain accurate LLM-generated code translations by leveraging pseudocode.
+Differently, we study the effectiveness of emulating semantic translation for code via pseudocode, where LLMs first interpret the semantics of original programs with pseudocode and then produce translated programs.
+We identify the complementary role of pseudocode-based translation for conventional direct translation, as well as its helpfulness in decoupling task burdens and bridging the differences between the features and APIs in varying PL pairs.
+The findings may guide practitioners to obtain accurate LLM-generated code translations in a new manner through abstraction.
 In addition, some works attempted to debug the erroneous code translations, e.g., by locating code snippets leading to inconsistent behaviors between the original and translated programs to guide repair (Wang et al., 2023b) and conducting certain types of repair based on fixed templates (Ramos et al., 2024).
-The methods mainly debug minor mistakes in the translated code.
+These methods mainly debug minor mistakes in translation results.
 They are generally orthogonal to our pseudocode-based code translation approach, which aims to reduce the occurrence of errors in the translation results and may also facilitate patch generation.
@@
 They also reveal that LLMs can generate programs in different PLs based on pseudocode and synthesize pseudocode for concrete programs, which motivates our study.
-In addition to code generation, pseudocode has also proven useful in cross-PL code retrieval by embedding structured pseudocode representations (Kulkarni et al., [n.
-d.]), in binary function similarity and vulnerability search by leveraging decompiled pseudocode to obtain platform-agnostic semantic representations (Zhang et al., 2022), and in generating descriptive summaries for stripped binaries using pseudocode extracted by decompilers as an expert-guided signal (Ye et al., 2023).
+In addition to code generation, pseudocode has also proven useful in cross-PL code retrieval by embedding structured pseudocode representations (Kulkarni et al., ), in binary function similarity and vulnerability search by leveraging decompiled pseudocode to obtain platform-agnostic semantic representations (Zhang et al., 2022), and in generating descriptive summaries for stripped binaries using pseudocode extracted by decompilers as an expert-guided signal (Ye et al., 2023).
 All these applications leveraged the advantage of pseudocode in succinctly capturing the intent and logic of programs to facilitate code comprehension and semantic representation (Oda et al., 2015).
-This work studies the helpfulness of pseudocode for another common software development task, i.e., code translation, where pseudocode serves as a semantic interpretation to facilitate semantic translation.
+This work studies the helpfulness of pseudocode for another common software development task, i.e., code translation, where pseudocode serves as a semantic interpretation for emulating semantic translation.
 The extensive investigation of pseudocode’s effectiveness in code translation yields various findings, which may guide practitioners to obtain accurate LLM-generated code translations based on pseudocode.
@@
 In this work, we empirically study the effectiveness of pseudocode-based code translation with LLMs, which explicitly emulates the human practice of semantic translation.
-By investigating the performance of five popular LLMs with pseudocode-based code translation on 9,690 translation tasks across six popular PLs, we reveal that pseudocode-based code translation can effectively complement the widely adopted direct code translation approach for various pairs of source and target PLs.
-We also demonstrate the effectiveness of pseudocode as a general and effective intermediary, as well as the further potential of translation based on higher-quality pseudocode and the bottleneck in both code understanding and generation.
-Our case studies further identify the concrete advantages and limitations of pseudocode-based code translation.
-Based on these findings, we suggest the hybrid use of pseudocode-based and direct code translation to enhance code translation accuracy, as well as discuss future research directions to further unleash the potential of pseudocode in code translation.
+By investigating the performance of five popular LLMs with pseudocode-based code translation on 9,690 translation tasks across six popular PLs, we reveal that pseudocode-based translation can effectively complement the widely adopted direct translation approach for various pairs of source and target PLs.
+We also identify the highly benefited PL pairs and demonstrate the effectiveness of pseudocode as a general and effective intermediary, as well as the further potential of translation based on higher-quality pseudocode and the bottleneck in both code understanding and generation.
+Our case studies further identify the advantages and limitations of pseudocode-based code translation.
+Based on these findings, we suggest the hybrid use of pseudocode-based and direct translation approaches to enhance code translation accuracy.
+We also discuss future research directions to further unleash the potential of pseudocode in code translation.
+Acknowledgments
+We would like to thank the anonymous reviewers for their insightful comments.
+This work is partially supported by the National Natural Science Foundation of China (Grant No.
+92582201), the Hong Kong SAR Research Grant Council/General Research Fund (Ref No.
+16210725), and the Hong Kong SAR Research Grant Council/Theme-based Research Scheme (Ref No.
+T41-517/25-N).